### v1

# Questions: A Taxonomy for Critical Reflection in Machine-Supported Decision-Making

Abstract.
Decision-makers run the risk of relying too much on machine recommendations. Explainable AI, a common strategy for calibrating reliance, has mixed and even negative effects, such as increasing overreliance. To cognitively engage the decision-maker and to facilitate a deliberate decision-making process, we propose a potential ‘reflection machine’ that supports critical reflection about the pending decision, including the machine recommendation. Reflection has been shown to improve critical thinking and reasoning, and thus decision-making. One way to stimulate reflection is to ask relevant questions. To systematically create questions, we present a question taxonomy inspired by Socratic questions and human-centred explainable AI. This taxonomy can contribute to the design of such a ‘reflection machine’ that asks decision-makers questions. Our work is part of the growing research on human-machine collaborations that goes beyond the paradigm of machine recommendations and explanations, and aims to enable greater human oversight as required by the European AI Act.
Reflective Practice, Socratic Questions, Cognitive Intervention, Human-AI Collaboration, Decision-Support Systems, Appropriate Reliance
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Conference; 978-1-4503-XXXX-X/18/06isbn: ;ccs: Human-centered computing Collaborative interactionccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Human-centered computing Computer supported cooperative work
1. Introduction
Reflective thinking is an important skill in professional practices (Schön, 1983). It is beneficial in, for example, management (Cunliffe, 2002), law (Casey, 2014), and healthcare (Mann et al., 2009; Sandars, 2009), and thus part of many educational curricula (Boud and Walker, 1998). Reflection, as John Loughran (1996, p14) states, is a “deliberate and purposeful act”, which, according to John Dewey (1933, p9), consists of “active, persistent and careful consideration”. During decision-making, reflection has the potential to improve reasoning, judgement, and problem-solving (Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Mamede and Schmidt, 2023), as it allows to evaluate the validity of information and assumptions (Mezirow, 1990). Consequently, reflection has been shown to improve strategic decisions (Walger et al., 2016) and diagnostic accuracy (Mamede et al., 2008; Prakash et al., 2019; Hess et al., 2015). In addition, reflection enables decision-makers to be more aware of the reasons for their decisions, which supports them in taking responsibility and being accountable. Ideally, responsible decisions lead to more desirable and fairer outcomes for all parties involved. In view of these effects, it seems sensible to encourage and support reflection during decision-making. Especially given the current paradigm in which many high-stakes decisions are supported by machine recommendations.
Decision support systems (DSS) assist decision-makers in domains like law, finance and healthcare. Physicians, for example, can use clinical decision support systems to find a treatment option for a patient. Studies show, however, that operators tend to rely too much on these systems and accept incorrect recommendations (Dratsch et al., 2023), which is known as overreliance (Passi and Vorvoreanu, 2022). A common approach to mitigating overreliance is explainable AI (XAI) (Longo et al., 2024), which focuses on providing information about which input was decisive for a particular output. Provided explanations are however often not taken into account (van der Waa et al., 2021), or they can reinforce the acceptance of any machine recommendation regardless of its quality or correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021). In light of this, and to reduce harmful consequences of overreliance, policymakers and laws such as the European AI Act require more significant, i.e., effective, human oversight over algorithmic decisions, enabling decision-makers to adequately scrutinise machine recommendations before making a decision (Pariliament, 2024).
In this paper, we propose a potential solution for mitigating overreliance, namely supporting the decision-maker in critically reflecting on the pending decision, including the machine recommendation. By critical reflection, we mean the practice of evaluating relevant information, without taking information at face value, and scrutinising (taken for granted) assumptions (Hong and Choi, 2011; Sengers et al., 2005). In relation to a DSS, this means that the decision-maker should remain critical of the output and data, as well as be open to alternative approaches. As mentioned, reflection can increase reasoning and allows to identify the reasons for a decision (Ghanizadeh, 2017). We therefore hypothesise that critical reflection can increase cognitive engagement of the decision-maker and thus help to reduce over-reliance on decision-support systems.
In order to elicit reflection, the decision-maker can be asked questions, particularly in the form of the Socratic questioning method. The Socratic method, named after the Greek philosopher Socrates, involves systematically asking questions in order to explore complex ideas, to uncover and scrutinise assumptions, and to gain a better understanding of the topic under investigation (Paul, 2019). Although the Socratic method is slowly gaining traction in the field of human-machine collaboration (Liu et al., 2024a, b; Lara, 2021; Ang et al., 2023), there is no systematic approach as to how this questioning method can be applied to the the context of machine-supported decision-making. We therefore provide a taxonomy of questions for critical reflection by synthesising and combining prior work, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revised version of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010). In doing so, this paper provides an answer to the question: How to create relevant questions that stimulate critical reflection in the context machine-supported decision-making? Although our question taxonomy is applicable to any decision-making domain, we primarily provide examples from the area of clinical decision-making, where decisions are often complex and have major implications.
This paper makes the following contributions:
  • We make the case for a potential ‘reflection machine’ that supports critical reflection through questions, by, among other things, repurposing explanations (section 2.2), which functions as a potential solution to reduce overreliance on DSS (section 2.3).
  • We discuss literature that shows that reflection increases cognitive engagement and reasoning, and thus decision-making, and that questions can help stimulate reflection (section 2.4), in particular in the form of Socratic questions (section 2.5).
  • We propose a new question taxonomy to help create relevant questions to stimulate reflection during machine-supported decision-making (section 4). In doing so, we concretise current conceptual proposals for human-machine collaborations that go beyond machine recommendations and explanations. This taxonomy provides designers and developers with the foundations for designing a ‘reflection machine’ that asks questions to the decision-maker.
2. Background
In the following section we will briefly discuss background work that led us to create our question taxonomy. In order to address the problem of overreliance on machine recommendations and explanations (section 2.1), we begin with a short review of the relevant literature on human-machine collaboration, particularly in relation to the human factor in explainable AI (section 2.2), and alternative strategies to reduce overreliance (section 2.3). From this, we derive that cognitive engagement is a promising approach. We identify the benefits of reflective thought for cognitive engagement by promoting critical thinking, reasoning, and thus responsible decision-making (section 2.4). Finally, we suggest that questions can be useful to stimulate reflection. We focus on Socratic questions, which are widely used in education (section 2.5). In short, to arrive at our taxonomy, we combine relevant insights from empirical findings and literature on human-machine collaboration, education, and philosophy.
2.1. Overreliance in Human-Machine Decision-Making
A meta-study of 106 experimental studies on human-machine collaboration found that decision-making performance generally decreases when decision-makers use decision-support systems (DSS) (Vaccaro et al., 2024). Decision-makers may tend to prefer machine recommendations over their own judgement (Tschandl et al., 2020), even when the recommendations are wrong (Jacobs et al., 2021; van der Stigchel et al., 2023). The acceptance of incorrect recommendations is referred to as overreliance (Passi and Vorvoreanu, 2022). As briefly mentioned in the introduction, a prominent method to calibrate reliance on DSS is explainable AI (XAI) (Vasconcelos et al., 2023). It is assumed that insight into how a DSS computed a decision will help the operator to assess the machine recommendation and thus make an informed decision about whether or not to follow the recommendation. The aforementioned meta-study on human-machine collaboration found that explanations do not lead to significant synergy effects in terms of the performance of the human-machine team (Vaccaro et al., 2024). On the contrary, explanations may increase the likelihood of operators accepting a machine recommendation regardless of its correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021), thereby increasing overreliance (Vered et al., 2023). One reason for this is that the focus on numerical explanations can give a false sense of objectivity, leading operators, regardless of their expertise, to overestimate the capabilities of a DSS (Ehsan et al., 2024). As such, the provision of additional information in the form of explanations does not automatically lead to more cognitive engagement of the decision-maker (Gajos and Mamykina, 2022).
2.2. Human-Centred XAI
Work is being done on more effective explanations by moving from the current algorithmic focus of XAI methods to a broader, socio-technical perspective of explanations (Liao and Varshney, 2022; Ehsan et al., 2021; Ehsan and Riedl, 2024; Ghassemi et al., 2021). Given the diverse needs of different stakeholders in the human-machine collaboration ecosystem, e.g., as debugging models, assessing regulatory compliance, making informed decisions, or contesting automated decision, current explainable AI methods that focus on providing a technical answer to how a DSS computed a particular output may not be always appropriate (Chari et al., 2024; Liao and Varshney, 2022; Brennen, 2020). As one participant working on clinical decision-support systems mentioned in a study by Liao et al. (2020) on human-centred explanations:
“[explanations by system rationale] are essentially ‘this is how I do it, take it or leave it’. But doctors don’t like this approach…Thinking that [AI is] giving treatment recommendations is the wrong place to start, because doctors know how to do it. It’s everything that happens around that decision they need help with… more discussions about the output, rather than how you get there” (p6).
So even if it is known how a DSS computed an output, operators might be more interested in contextual factors (Zednik, 2019; Ehsan and Riedl, 2024), which relate to questions like, “Is the data representative of the current situation?”, “Is the data up-to-date?”, “Are there any other overlooked possibilities?”. To stimulate “more discussions about the output” (study participant in Liao et al., 2020), we want to utilise current explainable AI methods and repurpose explanations to help formulate and inform questions that can be asked of decision-makers. As such, we consider “explanations as a means to help a data subject [and decision-maker] act rather than merely understand” (Wachter et al., 2017, p843), where ‘act’ in this case is deliberate decision-making.
2.3. Alternative Approaches to Reduce Overreliance
To address the remaining challenge of overreliance, scholars explore different interactions with machine recommendations and explanations that attempt to cognitively engage the decision-maker in the decision-making process. We will briefly mention three relevant approaches: 1) promoting cognitive engagement through interventions, 2) supporting the decision-maker to make their own decision by presenting evidence for and against, and 3) formulating explanations as questions.
First, cognitive interventions are based on the dual-process model of reasoning, and aim to interrupt the decision-maker’s habitual thinking (system 1) in order to encourage analytical thinking (system 2) (Lambe et al., 2016). Cognitive interventions can take the form of checklists, instructions for analytical thinking, or reflection, for example (Lambe et al., 2016; Prakash et al., 2019). In the context of human-machine collaboration, one study tested a cognitive intervention strategy by presenting explanations at different times, such as delayed or on-demand. These interventions were found to reduce, yet not eliminate, overreliance on DSS compared to the direct presentation of explanations (Buçinca et al., 2021).
Second, instead of presenting a recommendation and justifying it with an explanation that the decision-maker must accept or reject, a so-called hypothesis-driven recommender shows evidence for and against a decision (Miller, 2023). In this way, the system supports the decision-maker’s cognitive process and gives them the control to make an informed decision themselves. Another study found that while presenting evidence for and against did little to improve diagnostic accuracy, physicians valued the reflective aspect of it and had more confidence in their final decisions (Cabitza et al., 2023).
Third, in an experiment, causal explanations were presented both as statements and questions, and compared with each other (Danry et al., 2023). An explanation as a statement is, for example, “The Diagnosis $Y$ is because of symptom $x$”, and formulated as question: “If the patient has symptom $x$, does it necessarily mean they have condition $Y$?”. The study found that explanations framed as questions improved human judgement about the logical validity of the information provided. Similar to the hypothesis-driven recommender, questions help the decision-maker to think for themselves (Ryan and Deci, 2000), leading to more cognitive engagement (Gajos and Mamykina, 2022).
In addition to these approaches, it has been suggested that DSS should take on roles other than that of a recommender, such as an analyser presenting counterarguments, or as a devil’s advocate raising objections and challenging the decision-maker (Chiang et al., 2024; Sarkar, 2024; Ma et al., 2024a; Miller, 2023). Insights from decision-making in human groups show that these different roles, i.e., devil’s advocate and analyser, have positive effects on decision performance (Schwenk, 1984; Schweiger et al., 1989). Further, other authors argue that DSS should support deliberation and facilitate discussion between human and machine (Ma et al., 2024b; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023). Similarly, we imagine a ‘reflection machine’ that functions as a cognitive intervention by posing questions to the decision-maker to encourage them to reflect on the pending decision, including machine recommendations. Our question taxonomy for creating relevant, thought-provoking questions forms the foundation for the design of such a ‘reflection machine’.
2.4. Reflection and Responsible Decision-Making
Reflection proved to be the most effective form of cognitive intervention for increasing cognitive engagement, among other methods such as checklists or feedback (Lambe et al., 2016; Prakash et al., 2019). Reflection, also referred to as reflective practice (Schön, 1983), or reflective thought, is a multi-faceted concept with multiple meanings (Nguyen et al., 2014; Rogers, 2001; Mann et al., 2009). We adopt the definition of the American philosopher John Dewey (1933), who defines reflective thought as the “active, persistent and careful consideration of any belief or supposed form of knowledge in light of the grounds that support it and further conclusion to which it tends” (p9). Consequently, reflection can promote reasoning, critical thinking and problem-solving, and thus improve decision-making (Schmidt and Mamede, 2023; Mamede and Schmidt, 2023; Lambe et al., 2016; Prakash et al., 2019; Walger et al., 2016; Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Hess et al., 2015). While reflection can occur before, during or after an action (Rogers, 2001), we focus on reflection during decision-making, i.e., reflection-in-action (Schön, 1983).
In areas with far-reaching decisions, such as law or healthcare, decision-makers have a professional responsibility, such as physicians who are committed to the well-being of patients. This professional responsibility involves an epistemological responsibility, which means that decision-makers are responsible for gathering and evaluating relevant information, and knowing the reasons for a particular course of action (van Baalen and Boon, 2015). In regard to past actions, knowing and providing reasons for an action allows for backward-looking responsibility or accountability. More importantly, however, in regard to future actions, knowing the reasons allows to take forward-looking responsibility by determining and directing actions to achieve a desired state (Loughran, 1996). By making it possible to be aware of the reasons for a decision, to formulate and weigh them up accordingly, and to change them if necessary (Mezirow, 1990), reflection in individual cases, i.e., reflection-in-action, can increase epistemological responsibility, both in a backward-looking and, more importantly, in a forward-looking sense. Ideally, this forward-looking responsibility leads to fairer and better outcomes for all parties involved (Ueda et al., 2024). Reflection-in-action could, for example, prevent a physician from jumping to conclusions after seeing seemingly similar patient cases throughout the day, which could lead to tunnel vision (Mamede and Schmidt, 2023). Consequently, to cite Dewey again, reflection “converts action that is merely appetitive, blind and impulsive into intelligent action” (Dewey, 1933, p17).
2.5. The Socratic Questioning Method
A common technique for facilitating reflection is to ask questions (Osmond and Darlington, 2005). One questioning technique that is widely used in education is the Socratic method. The Socratic method consists of systematically asking questions that help to clarify concepts, improve understanding, and uncover gaps in knowledge (Paul, 2019). Further, it helps to discover own thoughts, to analyse assumptions, information and inferences, and to arrive at own judgements through own reasoning (Paul, 2019). Besides, answering questions promotes reasoning (Alfano, 2019). Socratic questions are thus closely linked to critical thinking, which aims at “judging in a reflective way what to do or what to believe” (Facione, 2000).
In general, independent thinking can promote the decision-autonomy of the decision-maker, which in turn can affect their motivation and well-being (Ryan and Deci, 2000). Findings from educational research suggest that motivation increases cognitive engagement (Rotgans and Schmidt, 2011; Blumenfeld et al., 2006) and even functions as a antecedent to it (Singh et al., 2022). As soon as students have the opportunity to participate in the organisation of learning activities, their interest and motivation increase (Blumenfeld et al., 2006). Without claiming a linear causality, questions seem to encourage reflection, which can increase decision-making autonomy, which can increase motivation, which can increase engagement.
In view of this, and the connection between critical reflection and reasoning and the associated positive effect on decision-making, we see potential in the transfer of questions, in particular through the Socratic method, from the education sector to the context of machine-supported decision-making. The Socratic method is slowly gaining traction in the field of human-machine collaboration, with a conceptual proposal of a virtual assistant that promotes reflection through Socratic questions (Lara and Deckers, 2020), and an educational chatbot based on a large language model and the Socratic method, which was found to improve critical reflection compared to traditional chatbots (Favero et al., 2024). There is, however, no systematic approach as to how the Socratic method and questions for critical reflection in general can be transferred to the area of machine-supported decision-making. This is partially because the Socratic method is primarily intended for the debate of philosophical issues. As such, a current taxonomy for Socratic questions (Paul, 2019, 2010) is primarily aimed at teachers and students to probe and foster thinking. We therefore aim to address this gap by transferring this taxonomy for Socratic questions and combining it with other work. In doing so, we concretise aforementioned conceptual proposals for human-machine collaborations that go beyond the traditional paradigm of machine recommendations and explanations (Ma et al., 2024b; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023).
3. Methodology
In the following, we will discuss how we have synthesised three earlier works in order to derive our question taxonomy, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revision of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010).
In the first step, we determine the elements for reflection, i.e., what the question shall relate to. To do so, we derive the overall structure of our taxonomy from a taxonomy for Socratic questions (Paul, 2019), which divides questions into the following components of reasoning:
  • Purpose: Questions that relate to the goal or objective.
  • Question at issue: Questions that relate to the problem or issue that gave rise to the question.
  • Information: Questions that relate to background information, such as data, facts, observations or experiences.
  • Interpretation and Inferences: Questions that relate to how meaning was derived and conclusions drawn.
  • Concepts: Questions that relate to underlying theories, definitions, or models that define a thought.
  • Assumptions: Questions that relate to taken for granted presuppositions of a thought.
  • Implications and Consequences: Questions that focus on implications.
  • Point of View: Questions that relate to the frame of reference or perspective of a thought.
We translate these categories and apply them to the context of decision-making in which operators use DSS. As such we the derive the following question categories: case information, e.g., input data, (information), relevance of data (information, interpretation and inference), dataset (concepts), causal structure of recommendation (interpretation and inferences), alternatives to recommendation (question, purpose), assumptions and expectations of decision-maker (assumptions, interpretation and inferences), stakeholder preferences (point of view), consequences of recommendation (implications), what-to-be that outcome (purpose), and model behaviour (assumptions) (see column ‘Element for Reflection’ in table 1).
In the second step, we identify elements, such as data or model behaviour through explanations, that can serve as information for the basis of the content of questions, i.e., how the questions can be enriched. To make appropriate use of the information derived from XAI, and given the various techniques available, we turn to a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022). The authors cluster more than 50 common questions operators have towards a DSS into the following categories:
  • How (global model-wide): Questions about the general logic of the model.
  • Why (a given prediction): Questions about the reasons for a prediction.
  • Why Not (a different prediction): Questions about the difference to an expected outcome.
  • How to be That (a different prediction): Questions about how to change
  • How to still be This (the current prediction): Questions about possible changes to still get the same outcome.
  • What if: Questions about changed outcome based on changed input.
  • Performance: Questions about the performance of the DSS.
  • Data: Questions about training data.
  • Output: Questions about how to apply or use the DSS output.
The authors use these question categories to map them to existing XAI techniques, such as feature importance or counterfactuals. This mapping is intended to help designers and developers select the appropriate XAI technique to provide the relevant answer, i.e., explanation.
In view of the overlap between the categorisation of Socratic questions and the common questions operators have, we use the mapping of questions to XAI techniques provided by the authors of (Liao et al., 2020, 2021; Liao and Varshney, 2022). Similarly, we assign XAI techniques to our identified question types in order to enrich possible questions with information (column ‘Useful Information’ in table 1). On the one hand, this means that (some) explanations can be (re-)formulated as questions (Danry et al., 2023), e.g., “Does outcome $Y$ follow from feature $x$?”. On the other hand, questions can relate to the decisive information extracted from explanations, e.g., “Is feature $x$ the relevant factor to focus on?”. So compared to the XAI question bank (Liao et al., 2020, 2021; Liao and Varshney, 2022), we reverse the process by repurposing explanations in order to ask the operators questions that stimulate critical reflection, instead of starting with questions from operators in order to arrive at possible explanations.
In the final step, and since questions can stimulate different thinking processes, we draw inspiration, albeit only marginally, from a revised version of Bloom’s taxonomy in order to capture the scope of questions (Krathwohl, 2002; Forehand, 2010; Anderson and Krathwohl, 2001). In Bloom’s taxonomy, questions are categorised hierarchically into six levels of cognitive processes, which are: remembering, understanding, applying, analysing, evaluating, and creating. The last three levels, analysing, evaluating, and creating, are also components of critical thinking (Paul, 2019). We thus hypothesise that in order to effectively stimulate critical reflection, questions should address these three higher-level cognitive abilities, where:
  • Analysing means to break down information into parts and determining how these parts are connected, and identifying which parts are relevant or irrelevant. Verbs that relate to this cognitive process are, compare, criticize, differentiate, discriminate, deconstruct, inspect. For example, “What evidence is there for $Y$?”.
  • Evaluating means to judge information and its appropriateness, and to detect inconsistencies. Verbs that relate to this cognitive process are, assess, support, defend, validate, evaluate. For example, “How to justify $Y$?”
  • Creating means compiling information in a new way, or considering alternative hypotheses. Verbs that relate to this cognitive process are, create, develop, formulate. For example, “What could be done to minimise $Y$?”
This distinction is meant to help formulate concrete questions for specific use cases.
4. Taxonomy of Questions for Critical Reflection
In order to help create fruitful questions for critical reflection, we present our taxonomy of question types (table 1). We refer to table 2 for some example questions from the medical domain, where a physician uses a DSS in order to find a diagnosis or treatment option for a patient. For this, a collection of general clinical questions served as input (Ely, 2000).
Table 1. Our question taxonomy, which supports the systematic creation of questions that stimulate critical reflection during decision-making. From a taxonomy of Socratic questions we take the elements of thought (Paul, 2019), indicated in italic, and translate them to the decision-making process between human and DSS. We provide helpful information that can enrich the questions. To map existing XAI techniques to question types, we utilise a question bank for human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022). IDs marked with an asterisk (*) indicate questions that address the level of creating, according to Bloom’s categorisation of questions for education (Krathwohl, 2002). We refer to table 2 for sample questions.
ID Element for Reflection (What?) Description (Why?) Useful Information (How?)
Q1 Case Information
Questions to help further assess, inspect, scrutinise, and contextualise data points, to ensure its quality and reliability. Input Data (e.g., tabular patient data)
Q2 Relevance of Data
(Information, Interpretation and Inference)
Questions to help evaluate data to ensure that it is valid by adequately supporting the hypothesis / recommendation. Derived explanations (XAI) can be framed as questions. Feature Contribution (e.g., SHAP (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016))
Q3 Dataset
(Concepts )
Questions to help inspect, assess and judge the assumptions built into the model to ensure that data adequately represents the phenomenon and that the conclusions derived are valid. Training Data, Datasheets (Gebru et al., 2021), Model Cards (Mitchell et al., 2019), FactSheets (Arnold et al., 2019), Global Feature Importance (Lundberg et al., 2020)
Q4 Causal Structure of Recommendation
(Interpretation and Inference)
Questions that help evaluate whether the outcome follows from data to ensure that the causal structure of the model / recommendation is sound. Can entail the previous questions of checking whether the data (Q1&Q2) and concept (Q3), i.e., model, are reliable and valid. Feature Contribution (e.g., (Lundberg and Lee, 2017), (Ribeiro et al., 2016))
Q5* Alternatives to Recommendation
(Question, Purpose)
Questions that help to consider other possibilities and create other options for actions, to ensure that (larger) solution space is considered. Contextual Information, Differential diagnosis, Patient Preferences (Q7)
Q6 Assumptions and Expectations of Decision-Maker
Questions that elicit taken for granted assumptions to ensure that the decision-maker is aware of their reasons and can direct their actions accordingly. General Questions
Q7* Stakeholder Preferences
(Point of View)
Questions that help the decision-maker enquire and take into account the preferences or needs of the people concerned (e.g., patient), in order to ensure a better or broader understanding of the problem at hand. General questions
Q8* Consequences of Recommendation
Questions that help elicit anticipation and forward-looking responsibility to ensure that (unintended) consequences are considered and mitigated. General questions
Q9* What to be that Outcome
Questions to support the exploration of alternative courses of action, by gaining insights from hypothetical scenarios derived from the DSS. Feature Perturbation, Counterfactuals (Karimi et al., 2021; Poyiadzi et al., 2020)
Q10 Model Behaviour
Questions that help to assess and evaluate the rules and thresholds of the model. In particular, areas of decision limits, i.e. until when does the result remain the same and when does it change. Allows to learn about causal structure (Q4). Decision Boundaries, Feature Perturbation, Counterfactuals, Model Cards (Mitchell et al., 2019)
Table 2. Based on our question taxonomy in table 1, we provide some sample questions that relate to a medical use case. A collection of general clinical questions serves as input (Ely, 2000). Relevant questions will vary from domain to domain, these questions should thus be taken as suggestions.
ID Element for Reflection Sample Questions
Q1 Case Information Is the provided data complete? How do you think the patient understood the self-assessment questionnaire? When was the indicated surgery performed? Why did provider $x$ treat the patient this way?
Q2 Relevance of Data What are the criteria for diagnosis of condition $y$? Could symptom $x$ be condition y or be a result of condition $y$? What is the likelihood that symptom $x$ is coming from condition $y$?
Q3 Dataset Does the data adequately represent the phenomena under investigation? Is the model up-to-date? Is there relevant data missing in the dataset? Did you consider $x$, which the model does not consider?
Q4 Causal Structure of Recommendation What is the likelihood that this patient has condition y (given findings $x_{1}$, $x_{2}$,…, $x_{n}$)? Which information supports/contradicts diagnosis? How good is test x in situation y?
Q5 Alternatives to Recommendation Are there previous patients with a similar profile who received a different treatment? Are there aspects that might have been overlooked? Is the alternative easy to reject? How do you distinguish between conditions $y_{1}$, $y_{2}$? Could this patient have condition y (given findings $x_{1}$, $x_{2}$)?
Q6 Assumptions and Expectations of Decision-Maker How does the machine recommendation compare to your assumptions? How aggressive/conservative should I be in situation $y$? What are you taking for granted? Are there alternative assumptions you could make?
Q7 Stakeholder Preferences Does the patient have any preferences that might require a procedure that differs from the recommendation? Does the patient have strong preferences for treatment $y$?
Q8 Consequences of Recommendation Are there any unintended consequences of treatment $y$? What are the ethical/legal considerations in situation $y$? What are the administrative considerations in situation y?
Q9 What to be that Outcome Is it possible to change the patient’s expectations to increase the likelihood of effective surgery?
Q10 Model Behaviour Would you suggest the same treatment if the patient were 5 years older?
Q1. Questions can address available information, such as data in the form of provided case information. Previous operations on the spine, for example, might be indicated by integers, but this value does not specify how long ago the surgery was performed and at which location of the spine it was carried out - both aspects that can influence the effectiveness of further operations. Questions can help the decision-maker to inspect and contextualise the data provided to give it semantic depth. In addition to ensuring the quality of the data, question can relate to the reliability of the data provided to ensure that inconsistencies, such as measurement errors, are ruled out, or that the data adequately reflect the intended purpose, such as the pain scores self-assessed by patients using a questionnaire.
Q2. The decisive factors for a specific machine recommendation, i.e., relevance of single data points, are another piece of information that can be questioned. XAI methods, such as SHAP (Lundberg and Lee, 2017) or LIME (Ribeiro et al., 2016), make it possible to extract the contribution of the various features to the outcome and rank them accordingly. In order to stimulate critical reflection, it is possible to use that information to generate an evaluative question, like “Is data point $x$ the one to focus on?” or “How relevant is data $x$?”. By questioning the relevance of decisive data points, the decision-maker is encouraged to contextualise the information, e.g., symptom, in a specific case. To do so, the decision-maker may wish to check and validate the machine recommendation by other means, e.g., by conducting additional tests. So while XAI can show which features have contributed to an outcome, domain knowledge is necessary to evaluate whether these features are important. Questioning the relevance of the most contributing data point also raises the question whether there are other factors that are overlooked in the decision. In addition, considering that different XAI techniques can result in different rankings of the contributing effects of the individual features (Saarela and Jauhiainen, 2021), it is important for decision-makers not to take the extracted information provided by XAI at face value. The contradiction between two opposing feature ranks by two different XAI techniques could also lead to a question like “Is $x$ or $z$ more relevant?”. Consequently, questions of this type (Q2) might overlap with Q4 that relates to the causal structure of recommendations.
Q3. Questions can address the concepts of an approach or problematisation, which in the context of joint decision-making translates to the dataset on which the model is built. Cardiovascular diseases, for example, show different symptoms in men and women. Most of the collected data that is used to train DSS, however, stems from male patients, which is why DSS often lead to misdiagnosis in women (Li et al., 2023). The problem in this case is that certain populations or phenomena are overrepresented in data, which leads to a skewed data distribution (Ueda et al., 2024). Besides, data is often only a partial representation of the phenomena under investigation (Broussard, 2019; Binns et al., 2018), as it is impossible to quantify every aspect of the complexity of social life (Birhane, 2021). Data might have been left out at training the model for different reasons, or was not available in the first place. Consequently, questions like “Does the data adequately represent the phenomena under investigation?”, or “Is more evidence necessary?”, can help to validate the functioning and appropriateness of the model and to learn about its limitations. In addition, questions like “Have you considered information x?” could be helpful if certain data is not included in the dataset and is therefore not considered by the model, although it is (now) known that the missing information contributes to the outcome. So compared to Q2, which is more concerned with the question of what information is relevant in an existing dataset, Q3 asks whether the dataset itself contains the relevant information. To create such questions, documentation of the dataset can be very useful (Gebru et al., 2021; Arnold et al., 2019).
Q4. To further assess the interpretation and inferences of the DSS, questions can help to investigate the (apparent) causal structure of provided case information and resulted machine recommendation. Again, based on feature contribution it is possible to ask whether and how the machine recommendation follows from data point $x$, for instance “Does diagnosis $Y$ follow from symptom $x$?”. So although this question type ‘merely’ reformulates a causal explanation, the aim is to help the decision-maker to analyse and evaluate information.
Q5. Questions can relate to the question or problem to which a DSS is trying to provide an answer, i.e., the purpose of the DSS. So similar to Q4 these questions can address the appropriateness of a recommendation. The focus of Q5, however, is on helping the decision-maker to consider alternatives to the recommendation. Different disease have common symptoms, so the physician wants to make sure that another possible diagnosis is not overlooked. For this, a differential diagnosis can be helpful (Mamede and Schmidt, 2023). At the same time, it can be possible to treat the same condition in different ways. Chronic lower back pain, for example, might be treated by surgical or non-surgical interventions (van Hooff et al., 2014), where the feasibility of each option can depend on constraints like available resources or the patient’s preferences (Q7). For example, if the purpose of a DSS is to optimise for full recovery, it might recommend surgery that involves a long stay in hospital; another purpose might be to promote the well-being of the patient, who may prefer to spend as little time as possible in hospital, so that the more suitable treatment could be a combined physical and psychological program. In this way, there is an overlap with question type Q7 that relates to stakeholder preferences. Nevertheless, the focus of Q5 is on whether alternative recommendations should be considered, and on supporting the decision-maker in developing alternative hypotheses and courses of action.
Q6. Questions can address assumptions and expectations of the decision-maker. On the one hand, these questions can help to elicit tacit knowledge and taken for granted assumptions, which might need to be re-evaluated on a case to case basis. This could prevent the decision-maker from jumping to conclusions. On the other hand, questions that address the expectations of the decision-maker can increase their confidence and allow them to arrive at their own decision. In other words, these questions can help to increase the decision-autonomy of the operator. Example questions could be, “Does the recommendation match your assumptions? If so, why (not)?”. General questions allow to take into account domain expertise that can address factors which might not be available in quantified data (Q3). As such, questions could allow the decision-maker to obtain a more holistic picture of the problem that goes beyond static rules implemented in DSS (van Baalen and Boon, 2015) (Q5).
Q7. Questions can relate to different perspectives or points of view. Especially in decision context in which decision-maker and decision-subject are different persons, such as physician and patient, it is important to take into account stakeholder preferences. As such, questions like “Have you considered the preferences of the patient?”, or “Are there certain situational circumstances that prevent the patient from recovering from surgery for several weeks?”, can help to ensure the decision-maker considers other parties involved and treat the patient in a more personalised way. Ideally, enquiring about the patient’s preferences is part of the physician’s routine practice during consultation, yet these questions can function as a form of safeguard.
Q8. Questions can help to anticipate (unintended) consequences of a decision. These questions can elicit forward-looking responsibility of the decision-maker, for example, by asking “Have you considered the implications of administering drug $x$?” While the decision-maker has ideally weighed up the implications, these questions can function as a reminder. Furthermore, with insights from the dataset (Q3) attention could be drawn to certain limitations of a DSS. Compared to Q3, the focus of Q8 is on assessing the impact of a recommendation that may be based on underrepresented data, and supporting the decision-maker in considering appropriate mitigation strategies.
Q9. Questions can take the form of what-to-be that outcome in order to further investigate the space of different alternatives. By creating hypothetical scenarios, the feasibility of desired outcomes can be investigated. So for example, ‘Is it possible to lower level $z$, this could probably make treatment $y$ more effective?” In doing so, the decision-maker is encouraged to consider prior and smaller interventions that could increase the expected effectiveness of a desired treatment. To create such question, feature perturbation and counterfactual explanations are helpful (Karimi et al., 2021; Poyiadzi et al., 2020). In contrast to counterfactual explanations, which provide information about what the outcome would have been if the data had been like this, these questions are about whether it is possible to change certain data points in order to make a particular outcome more likely. For example, a counterfactual explanation would be “If finding $x$ were not present, the condition would be $y$”, whereas the question turns into “Is it possible to lower $x$, so that the likelihood of condition $y$ increases?”. Although these questions ideally lead to more actionable insights, they might imply unfeasible or unattainable changes and actions, e.g., changing the age of a patient. These hypothetical scenarios can nevertheless provide insights into inbuilt associations and thresholds of the DSS decision-boundaries.
Q10. Questions can address assumptions, as mentions before. In addition to the assumptions of the decision-maker, questions can refer to built-in assumptions or rules of model behaviour. For example, if a DSS makes a decision based on age, with a threshold of 50 years, the decision-maker may want to re-evaluate the decision, if the patient is close to that threshold, e.g., 48 years old. An example question could be “Would you recommend the same treatment if the patient were 3 years older, which might decrease its effectiveness?”. These questions might also relate to the relevance of data points (Q1), as single data points often function as proxies for other characteristics. Age, for example, might be an indication for how well the patient will recover from surgery, yet physical fitness might also play a role. This means, focusing only on single data points could be misleading. As such, these questions can also give the decision-maker a better understanding of the causal structure built into the model (Q4). Documentation on model behaviour can be a source from which questions can be derived (Mitchell et al., 2019).
5. Discussion, Future Work, and Concluding Remarks
We have proposed a taxonomy of question types to support the systematic creation of questions that can stimulate critical reflection during machine-supported decision-making. In doing so, we have focused on the medical domain by providing some example questions. We maintain that our taxonomy is transferable to other decision-making domains. Whilst the wording of the questions may change when applied to other contexts, the general structure and elements of critical reflection remain the same. Future work is required to test and validate the generalisability of our taxonomy. Next to that some considerations remain.
The tone of questions is important. For natural language explanations, it was found that the tone and assertiveness has an effect on decision-making (Calisto et al., 2023). Although questions are in general less assertive compared to explanations, it could be assumed that the tone of questions has a similar effect on how or whether the decision-maker engages with the questions. A question could be framed in a more imposing way, such as, “Can you check for symptom $x$?”, rather than in a suggestive way, e.g., “Have you checked for other symptoms?”. To examine the right tone, we have presented Bloom’s taxonomy. We argue that questions should follow the higher levels of critical thinking, that is evaluating, analysing, and creating. As such, it is likely to assume that the questions should be phrased in an open-ended manner. Simple yes-or-no questions could be overlooked or act as checkboxes without eliciting cognitive engagement. Besides, it is not necessarily the case that definite answers have to be found to questions. Rather, the posed question can lead to further questions and thus inquiry of the decision-maker (Paul, 2019). In order to support reflection effectively, it is therefore necessary to create questions in a controlled and systematic way. Consequently, it might be advisable to (currently) refrain from using large language models (LLMs), as the output risks being irrelevant and unreliable (Weidinger et al., 2022). Besides, conversational presentation of information in the form of chatbots might lead to further overreliance (Anderl et al., 2024), as humanised or anthropomorphised chatbots can lead to an illusion of reciprocity and care (Maeda and Quan-Haase, 2024). Future work could nevertheless investigate how to stimulate continuous reflection by processing the answers of the operator and asking follow-up questions. Ideally, an interactive interface would allow the decision-maker to modify data points accordingly. At the same time, however, decision-makers should not be burdened with too many or too frequent questions. In the aforementioned study on cognitive interventions, the cognitive load of these interventions on the decision-maker was reported, which is why participants favoured them less (Buçinca et al., 2021). The right balance of attention and engagement must be found. The operator should be enabled to constructively take the posed question into account, which also depends at least in part on the competences and confidence of the decision-maker. Ultimately, the cognitive burden is unavoidable, and inherent to critical reflection. As John Dewey (1933) mentions on the difficulty of reflective thinking:
yet thinking need not be reflective. For the person may not be sufficiently critical about the ideas that occur to him. He may jump at a conclusion without weighing the grounds on which it rests; he may forego or unduly shorten the act of hunting, inquiring; he may take the first ‘answer’ or solution, that comes to him because of mental sloth, torpor, impatience to get something settled. One can think reflectively only when one is willing to endure suspense and to undergo the trouble of searching (p17).
With this in mind, and given that questions during the decision-making process are likely to reduce efficiency by interrupting the workflow and taking up more time, the design and implementation of a potential ‘reflection machine’ that supports reflection seems less desirable. However, in domains with high-impact decisions, such as healthcare, it should be justified to make a deliberate and thus perhaps more time-consuming decision, as the outcome can have life-changing effects. Considering that DSS recycle past assumptions with potentially harmful consequences (Burrell, 2024), critical reflection becomes increasingly important to promote responsible decision-making. In particular, forward-looking responsibility, which can be strengthened through reflection, ideally leads to a more open and just future.
In this paper we provided a taxonomy of questions aimed at promoting critical reflection on machine recommendations. Careful selection and presentation of these question types in the right circumstances (e.g., when time for reflection is available), and appropriate combination with recommendations and explanations can support the effective human oversight required by the European AI Act.

### v2

@@
-Questions: A Taxonomy for Critical Reflection in Machine-Supported Decision-Making
+A Taxonomy of Questions for Critical Reflection in Machine-Assisted Decision-Making
 Abstract.
-Decision-makers run the risk of relying too much on machine recommendations.
-Explainable AI, a common strategy for calibrating reliance, has mixed and even negative effects, such as increasing overreliance.
-To cognitively engage the decision-maker and to facilitate a deliberate decision-making process, we propose a potential ‘reflection machine’ that supports critical reflection about the pending decision, including the machine recommendation.
-Reflection has been shown to improve critical thinking and reasoning, and thus decision-making.
-One way to stimulate reflection is to ask relevant questions.
-To systematically create questions, we present a question taxonomy inspired by Socratic questions and human-centred explainable AI.
-This taxonomy can contribute to the design of such a ‘reflection machine’ that asks decision-makers questions.
-Our work is part of the growing research on human-machine collaborations that goes beyond the paradigm of machine recommendations and explanations, and aims to enable greater human oversight as required by the European AI Act.
-Reflective Practice, Socratic Questions, Cognitive Intervention, Human-AI Collaboration, Decision-Support Systems, Appropriate Reliance
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Conference; 978-1-4503-XXXX-X/18/06isbn: ;ccs: Human-centered computing Collaborative interactionccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Human-centered computing Computer supported cooperative work
+Decision-makers run the risk of relying too much on machine recommendations, which is associated with lower cognitive engagement.
+Reflection has been shown to increase cognitive engagement and improve critical thinking and reasoning and therefore decision-making.
+However, there is currently no approach to support reflection in machine-assisted decision-making.
+We therefore present a taxonomy that serves to systematically create questions related to machine-assisted decision-making that promote reflection and thus cognitive engagement and ultimately a deliberate decision-making process.
+Our taxonomy builds on a taxonomy of Socratic questions and a question bank for human-centred explainable AI (XAI), and illustrates how XAI techniques can be utilised and repurposed to formulate questions.
+As a use case, we focus on clinical decision-making.
+An evaluation in education confirms the applicability and expected benefits of our taxonomy.
+Our work contributes to the growing research on human-AI interaction that goes beyond the paradigm of machine recommendations and explanations and aims to enable effective human oversight as required by the European AI Act.
+Reflection, Cognitive Engagement, Decision-Support Systems, Overreliance, Critical Thinking, Socratic Questioning, Human-AI Interaction
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; isbn: ;ccs: Human-centered computing Collaborative interactionccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Human-centered computing Computer supported cooperative work
 1. Introduction
-Reflective thinking is an important skill in professional practices (Schön, 1983).
-It is beneficial in, for example, management (Cunliffe, 2002), law (Casey, 2014), and healthcare (Mann et al., 2009; Sandars, 2009), and thus part of many educational curricula (Boud and Walker, 1998).
-Reflection, as John Loughran (1996, p14) states, is a “deliberate and purposeful act”, which, according to John Dewey (1933, p9), consists of “active, persistent and careful consideration”.
-During decision-making, reflection has the potential to improve reasoning, judgement, and problem-solving (Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Mamede and Schmidt, 2023), as it allows to evaluate the validity of information and assumptions (Mezirow, 1990).
+Reflection, also referred to as reflective practice or reflective thought, is an important skill in professional practices (Schön, 1983).
+It is beneficial in, for example, management (Cunliffe, 2002), law (Casey, 2014), and healthcare (Mann et al., 2009; Sandars, 2009), and is part of many educational curricula (Boud and Walker, 1998).
+As the American philosopher John Dewey (1933, p.9) states, reflection consists of an
+“active, persistent and careful consideration of any belief or supposed form of knowledge in light of the grounds that support it and further conclusion to which it tends”.
+As reflection helps to scrutinise the validity of information and assumptions (Mezirow, 1990), it has the potential to improve reasoning, judgement, and problem-solving (Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Mamede and Schmidt, 2023).
 Consequently, reflection has been shown to improve strategic decisions (Walger et al., 2016) and diagnostic accuracy (Mamede et al., 2008; Prakash et al., 2019; Hess et al., 2015).
-In addition, reflection enables decision-makers to be more aware of the reasons for their decisions, which supports them in taking responsibility and being accountable.
-Ideally, responsible decisions lead to more desirable and fairer outcomes for all parties involved.
-In view of these effects, it seems sensible to encourage and support reflection during decision-making.
-Especially given the current paradigm in which many high-stakes decisions are supported by machine recommendations.
-Decision support systems (DSS) assist decision-makers in domains like law, finance and healthcare.
-Physicians, for example, can use clinical decision support systems to find a treatment option for a patient.
+In addition, reflection enables decision-makers to recognise and articulate the reasons for their decisions, which helps them to direct their actions accordingly and take responsibility (forward-looking) and accountability (backward-looking).
+Responsible decision-making ideally leads to more desirable and fairer outcomes for all parties involved.
+In view of these effects, it seems sensible to encourage and support reflection during decision-making, especially as many high-impact decisions nowadays are informed by machine recommendations.
+Decision-support systems (DSS) assist decision-makers in domains like law, finance and healthcare.
+Physicians, for example, can use clinical DSSs to make a diagnosis or find a treatment option for a patient.
 Studies show, however, that operators tend to rely too much on these systems and accept incorrect recommendations (Dratsch et al., 2023), which is known as overreliance (Passi and Vorvoreanu, 2022).
-A common approach to mitigating overreliance is explainable AI (XAI) (Longo et al., 2024), which focuses on providing information about which input was decisive for a particular output.
-Provided explanations are however often not taken into account (van der Waa et al., 2021), or they can reinforce the acceptance of any machine recommendation regardless of its quality or correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021).
-In light of this, and to reduce harmful consequences of overreliance, policymakers and laws such as the European AI Act require more significant, i.e., effective, human oversight over algorithmic decisions, enabling decision-makers to adequately scrutinise machine recommendations before making a decision (Pariliament, 2024).
-In this paper, we propose a potential solution for mitigating overreliance, namely supporting the decision-maker in critically reflecting on the pending decision, including the machine recommendation.
-By critical reflection, we mean the practice of evaluating relevant information, without taking information at face value, and scrutinising (taken for granted) assumptions (Hong and Choi, 2011; Sengers et al., 2005).
-In relation to a DSS, this means that the decision-maker should remain critical of the output and data, as well as be open to alternative approaches.
-As mentioned, reflection can increase reasoning and allows to identify the reasons for a decision (Ghanizadeh, 2017).
-We therefore hypothesise that critical reflection can increase cognitive engagement of the decision-maker and thus help to reduce over-reliance on decision-support systems.
-In order to elicit reflection, the decision-maker can be asked questions, particularly in the form of the Socratic questioning method.
-The Socratic method, named after the Greek philosopher Socrates, involves systematically asking questions in order to explore complex ideas, to uncover and scrutinise assumptions, and to gain a better understanding of the topic under investigation (Paul, 2019).
-Although the Socratic method is slowly gaining traction in the field of human-machine collaboration (Liu et al., 2024a, b; Lara, 2021; Ang et al., 2023), there is no systematic approach as to how this questioning method can be applied to the the context of machine-supported decision-making.
-We therefore provide a taxonomy of questions for critical reflection by synthesising and combining prior work, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revised version of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010).
-In doing so, this paper provides an answer to the question: How to create relevant questions that stimulate critical reflection in the context machine-supported decision-making?
-Although our question taxonomy is applicable to any decision-making domain, we primarily provide examples from the area of clinical decision-making, where decisions are often complex and have major implications.
+Moreover, explainable AI (XAI) (Longo et al., 2024), a common approach for calibrating reliance, has mixed and even negative effects, such as increasing overreliance (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021).
+In light of this, and to reduce harmful consequences of overreliance, the European AI Act (European Parliament and Council of the European Union, 2024, §14) requires:
+[…] that natural persons to whom human oversight is assigned are enabled, as appropriate and proportionate: […] (b) to remain aware of the possible tendency of automatically relying or over-relying on the output produced by a high-risk AI system (automation bias), in particular for high-risk AI systems used to provide information or recommendations for decisions to be taken by natural persons;
+In this paper, we propose a potential solution for mitigating overreliance, namely cognitive engagement of the decision-maker through questions that encourage critical reflection on the decision at hand, including the machine recommendation and the data that led to it.
+By critical reflection, we mean the above-mentioned practice of evaluating relevant information, without taking information at face value, and analysing assumptions that may have been taken for granted (Hong and Choi, 2011; Sengers et al., 2005).
+In order to elicit reflection, the decision-maker can be asked questions, for which the Socratic questioning method offers a helpful starting point.
+The Socratic method, named after the Greek philosopher Socrates, involves systematically asking questions to explore complex ideas, uncover assumptions, and gain a better understanding of the topic under investigation (Paul, 2019).
+Although the Socratic method is slowly gaining traction in the field of human-machine interaction (Liu et al., 2024; Lara, 2021; Ang et al., 2023), there is as yet no systematic approach to applying this questioning method to the context of machine-assisted decision-making.
+To address this gap, we therefore provide a taxonomy of questions that serves the creation of questions for critical reflection in machine-assisted decision-making.
+For our taxonomy we synthesise prior work, namely 1) a taxonomy of Socratic questions (primarily for education) (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revised version of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010).
+We primarily focus on clinical decision-making, where a physician uses a DSS to find a treatment option for patients.
+In this context, there may be multiple solutions with uncertain outcomes, so deliberate decision-making is crucial.
+An evaluation of the applicability of our taxonomy in the education domain shows however that our taxonomy can be applied to other domains without adaptation.
+Overall, our proposed taxonomy builds on evidence from literature that critical reflection can increase cognitive engagement of the decision-maker.
 This paper makes the following contributions:
-  • We make the case for a potential ‘reflection machine’ that supports critical reflection through questions, by, among other things, repurposing explanations (section 2.2), which functions as a potential solution to reduce overreliance on DSS (section 2.3).
-  • We discuss literature that shows that reflection increases cognitive engagement and reasoning, and thus decision-making, and that questions can help stimulate reflection (section 2.4), in particular in the form of Socratic questions (section 2.5).
-  • We propose a new question taxonomy to help create relevant questions to stimulate reflection during machine-supported decision-making (section 4).
-In doing so, we concretise current conceptual proposals for human-machine collaborations that go beyond machine recommendations and explanations.
-This taxonomy provides designers and developers with the foundations for designing a ‘reflection machine’ that asks questions to the decision-maker.
+  • With respect to overreliance (section 2.1) and human-centred XAI (section 2.2), we add to the growing body of literature on approaches to increase cognitive engagement (section 2.3) and, in particular, work towards and concretise existing proposals to stimulate reflection.
+  • We discuss literature that shows that reflection increases cognitive engagement and reasoning, and thus decision-making (section 2.4), and that questions can help stimulate reflection (section 2.5).
+  • We present a new taxonomy to help formulate relevant questions that stimulate reflection in machine-assisted decision-making (section 4).
+This taxonomy supports the design and development of interventions that go beyond machine recommendations and explanations.
+In addition, our taxonomy can benefit other stakeholders such as developers, auditors, or consumers who need to remain critical in the development, evaluation, and use of DSS (section 6).
 2. Background
-In the following section we will briefly discuss background work that led us to create our question taxonomy.
-In order to address the problem of overreliance on machine recommendations and explanations (section 2.1), we begin with a short review of the relevant literature on human-machine collaboration, particularly in relation to the human factor in explainable AI (section 2.2), and alternative strategies to reduce overreliance (section 2.3).
-From this, we derive that cognitive engagement is a promising approach.
-We identify the benefits of reflective thought for cognitive engagement by promoting critical thinking, reasoning, and thus responsible decision-making (section 2.4).
-Finally, we suggest that questions can be useful to stimulate reflection.
-We focus on Socratic questions, which are widely used in education (section 2.5).
-In short, to arrive at our taxonomy, we combine relevant insights from empirical findings and literature on human-machine collaboration, education, and philosophy.
-2.1. Overreliance in Human-Machine Decision-Making
-A meta-study of 106 experimental studies on human-machine collaboration found that decision-making performance generally decreases when decision-makers use decision-support systems (DSS) (Vaccaro et al., 2024).
+In the following, we will discuss background work that led us to the creation of our question taxonomy.
+In order to sketch the problem of overreliance on machine recommendations and explanations (section 2.1), we begin with a short review of the relevant literature on human-machine interaction, particularly in relation to the human factor in explainable AI (section 2.2).
+From alternative interaction approaches, we derive that encouraging cognitive engagement is a promising method to reduce overreliance (section 2.3).
+We identify the benefits of reflective thought for cognitive engagement as it promotes critical thinking, reasoning, and thus responsible decision-making (section 2.4).
+Finally, we suggest that questions can be useful to stimulate reflection and thus cognitive engagement, and introduce the Socratic questioning method, which is widely used in education (section 2.5).
+In short, to arrive at our taxonomy, we combine relevant insights from empirical findings and literature on human-machine interaction, education, and philosophy.
+2.1. Overreliance on DSS
+Decision-support systems (DSS) generally process information to arrive at an output, i.e., recommendation, by using (simple) mathematical models, or more advanced machine learning models, including deep learning approaches.
+A meta-study of 106 experimental studies on machine-assisted decision-making found that decision-making accuracy generally decreases when decision-makers use DSS (Vaccaro et al., 2024).
 Decision-makers may tend to prefer machine recommendations over their own judgement (Tschandl et al., 2020), even when the recommendations are wrong (Jacobs et al., 2021; van der Stigchel et al., 2023).
-The acceptance of incorrect recommendations is referred to as overreliance (Passi and Vorvoreanu, 2022).
-As briefly mentioned in the introduction, a prominent method to calibrate reliance on DSS is explainable AI (XAI) (Vasconcelos et al., 2023).
-It is assumed that insight into how a DSS computed a decision will help the operator to assess the machine recommendation and thus make an informed decision about whether or not to follow the recommendation.
-The aforementioned meta-study on human-machine collaboration found that explanations do not lead to significant synergy effects in terms of the performance of the human-machine team (Vaccaro et al., 2024).
+The acceptance of incorrect recommendations is referred to as overreliance, which can be because of low vigilance or lack of cognitive engagement (Passi and Vorvoreanu, 2022).
+A common approach to mitigating overreliance is explainable AI (XAI) (Vasconcelos et al., 2023), which can, for example, provide information about which input was decisive for a particular output, i.e., recommendation.
+It is assumed that these explanations will help the operator to assess the machine recommendation and thus make an informed decision about whether or not to follow the recommendation.
+Provided explanations are, however, not always taken into account (van der Waa et al., 2021).
+Further, the aforementioned meta-study also found that explanations do not lead to significant synergy effects in terms of the performance of the human-machine team (Vaccaro et al., 2024).
 On the contrary, explanations may increase the likelihood of operators accepting a machine recommendation regardless of its correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021), thereby increasing overreliance (Vered et al., 2023).
 One reason for this is that the focus on numerical explanations can give a false sense of objectivity, leading operators, regardless of their expertise, to overestimate the capabilities of a DSS (Ehsan et al., 2024).
-As such, the provision of additional information in the form of explanations does not automatically lead to more cognitive engagement of the decision-maker (Gajos and Mamykina, 2022).
-2.2. Human-Centred XAI
-Work is being done on more effective explanations by moving from the current algorithmic focus of XAI methods to a broader, socio-technical perspective of explanations (Liao and Varshney, 2022; Ehsan et al., 2021; Ehsan and Riedl, 2024; Ghassemi et al., 2021).
-Given the diverse needs of different stakeholders in the human-machine collaboration ecosystem, e.g., as debugging models, assessing regulatory compliance, making informed decisions, or contesting automated decision, current explainable AI methods that focus on providing a technical answer to how a DSS computed a particular output may not be always appropriate (Chari et al., 2024; Liao and Varshney, 2022; Brennen, 2020).
-As one participant working on clinical decision-support systems mentioned in a study by Liao et al.
-(2020) on human-centred explanations:
+Consequently, the provision of additional information in the form of explanations does not automatically lead to more cognitive engagement of the decision-maker and more accurate decisions (Gajos and Mamykina, 2022).
+2.2. A Question Bank for XAI
+To provide more actionable explanations based on user needs, a question bank for explainable AI helps designers and developers select the appropriate XAI technique to provide a relevant explanation that supports the operator’s goal, e.g., informed decision-making.
+To this end, the question bank maps prototypical user questions to available XAI techniques (Liao et al., 2020, 2021; Liao and Varshney, 2022).
+Feature importance, for example, can provide insights into why-questions, such as “Why was this prediction given?”.
+Answering the (technical) question of how a DSS arrived at a particular output is however not always desired or helpful.
+As one participant working on clinical decision-support systems mentioned in the study on the XAI question bank:
 “[explanations by system rationale] are essentially ‘this is how I do it, take it or leave it’.
 But doctors don’t like this approach…Thinking that [AI is] giving treatment recommendations is the wrong place to start, because doctors know how to do it.
-It’s everything that happens around that decision they need help with… more discussions about the output, rather than how you get there” (p6).
-So even if it is known how a DSS computed an output, operators might be more interested in contextual factors (Zednik, 2019; Ehsan and Riedl, 2024), which relate to questions like, “Is the data representative of the current situation?”, “Is the data up-to-date?”, “Are there any other overlooked possibilities?”.
-To stimulate “more discussions about the output” (study participant in Liao et al., 2020), we want to utilise current explainable AI methods and repurpose explanations to help formulate and inform questions that can be asked of decision-makers.
-As such, we consider “explanations as a means to help a data subject [and decision-maker] act rather than merely understand” (Wachter et al., 2017, p843), where ‘act’ in this case is deliberate decision-making.
-2.3. Alternative Approaches to Reduce Overreliance
-To address the remaining challenge of overreliance, scholars explore different interactions with machine recommendations and explanations that attempt to cognitively engage the decision-maker in the decision-making process.
-We will briefly mention three relevant approaches: 1) promoting cognitive engagement through interventions, 2) supporting the decision-maker to make their own decision by presenting evidence for and against, and 3) formulating explanations as questions.
+It’s everything that happens around that decision they need help with… more discussions about the output, rather than how you get there” (Liao et al., 2020, p.6).
+While the XAI question bank allows for the design of more human-centred XAI, it is still very algorithmic- and data-centred, as we will show (Figure 1) and as the authors themselves note.
+Even if it is known how a DSS arrived at an output or recommendation, operators might be more interested in factors outside the model (Zednik, 2019; Ehsan and Riedl, 2024), which could relate to questions like, “What is missing from the dataset?”, “Are there any other overlooked alternative solutions?”, or “How to integrate this recommendation into the decision-process?”.
+Scholars therefore advocate moving from the current algorithmic focus of XAI methods to a broader, socio-technical perspective of explanations (Ehsan et al., 2021; Ehsan and Riedl, 2024; Ghassemi et al., 2021).
+With our proposed taxonomy, we want to stimulate “more discussions about the output” (study participant in Liao et al., 2020) by identifying and categorising relevant question types.
+As we will show, current XAI techniques can be utilised and explanations can be repurposed to formulate and inform possible questions.
+Framing causal explanations as questions, for example, was found to improve human judgement (Danry et al., 2023).
+As such, in line with actionability, we consider “explanations as a means to help a data subject [and decision-maker] act rather than merely understand” (Wachter et al., 2017, p.843), where ‘act’ in this case is deliberate decision-making through reflection.
+2.3. Increasing Cognitive Engagement
+To mitigate overreliance, scholars explore different interactions with machine recommendations and explanations that attempt to cognitively engage the decision-maker in the decision-making process.
+We will briefly mention four relevant approaches: 1) promoting cognitive engagement through interventions, 2) supporting the decision-maker to make their own decision by presenting evidence for and against, 3) formulating explanations as questions, and 4) proposals to stimulate reflection.
 First, cognitive interventions are based on the dual-process model of reasoning, and aim to interrupt the decision-maker’s habitual thinking (system 1) in order to encourage analytical thinking (system 2) (Lambe et al., 2016).
 Cognitive interventions can take the form of checklists, instructions for analytical thinking, or reflection, for example (Lambe et al., 2016; Prakash et al., 2019).
-In the context of human-machine collaboration, one study tested a cognitive intervention strategy by presenting explanations at different times, such as delayed or on-demand.
-These interventions were found to reduce, yet not eliminate, overreliance on DSS compared to the direct presentation of explanations (Buçinca et al., 2021).
+In the context of human-machine interaction, one study found that such interventions, i.e., presenting explanations delayed, on-demand, or only after the operator made a decision, can reduce, yet not eliminate, overreliance on DSS (Buçinca et al., 2021).
 Second, instead of presenting a recommendation and justifying it with an explanation that the decision-maker must accept or reject, a so-called hypothesis-driven recommender shows evidence for and against a decision (Miller, 2023).
-In this way, the system supports the decision-maker’s cognitive process and gives them the control to make an informed decision themselves.
-Another study found that while presenting evidence for and against did little to improve diagnostic accuracy, physicians valued the reflective aspect of it and had more confidence in their final decisions (Cabitza et al., 2023).
-Third, in an experiment, causal explanations were presented both as statements and questions, and compared with each other (Danry et al., 2023).
-An explanation as a statement is, for example, “The Diagnosis $Y$ is because of symptom $x$”, and formulated as question: “If the patient has symptom $x$, does it necessarily mean they have condition $Y$?”.
+In this way, the system supports the decision-maker’s reasoning process and gives them the control to make an informed decision themselves.
+A study found that while presenting evidence for and against did little to improve diagnostic accuracy, physicians valued the reflective aspect and had more confidence in their final decisions (Cabitza et al., 2023).
+Third, in an experiment causal explanations were presented both as statements and questions (Danry et al., 2023).
+For example, the explanation as a statement: “The patient has the flu because of a headache”, is formulated as a question: “If the patient has a headache, does it follow that they have the flu?”.
 The study found that explanations framed as questions improved human judgement about the logical validity of the information provided.
-Similar to the hypothesis-driven recommender, questions help the decision-maker to think for themselves (Ryan and Deci, 2000), leading to more cognitive engagement (Gajos and Mamykina, 2022).
-In addition to these approaches, it has been suggested that DSS should take on roles other than that of a recommender, such as an analyser presenting counterarguments, or as a devil’s advocate raising objections and challenging the decision-maker (Chiang et al., 2024; Sarkar, 2024; Ma et al., 2024a; Miller, 2023).
-Insights from decision-making in human groups show that these different roles, i.e., devil’s advocate and analyser, have positive effects on decision performance (Schwenk, 1984; Schweiger et al., 1989).
-Further, other authors argue that DSS should support deliberation and facilitate discussion between human and machine (Ma et al., 2024b; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023).
-Similarly, we imagine a ‘reflection machine’ that functions as a cognitive intervention by posing questions to the decision-maker to encourage them to reflect on the pending decision, including machine recommendations.
-Our question taxonomy for creating relevant, thought-provoking questions forms the foundation for the design of such a ‘reflection machine’.
+Similar to the hypothesis-driven recommender, questions help the decision-maker to think for themselves (Ryan and Deci, 2000), which leads to more cognitive engagement (Gajos and Mamykina, 2022).
+Fourth, it has been suggested that DSS should take on roles other than that of a recommender, such as a devil’s advocate raising objections and challenging the decision-maker (Chiang et al., 2024; Sarkar, 2024; Ma et al., 2024b).
+Insights from decision-making in human groups show that a devil’s advocate has positive effects on decision performance (Schwenk, 1984; Schweiger et al., 1989).
+Similarly, others argue that DSS should support deliberation and facilitate reflection (Ma et al., 2024a; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023; Cornelissen et al., 2022).
+In view of the above, we understand questions as cognitive interventions (1), which facilitate own reasoning and decision-making (2&3).
+With our taxonomy, we therefore aim to both build on and work towards these four approaches by offering more potential questions compared to 3) and concretising the proposals for promoting deliberation and reflection (4).
 2.4. Reflection and Responsible Decision-Making
-Reflection proved to be the most effective form of cognitive intervention for increasing cognitive engagement, among other methods such as checklists or feedback (Lambe et al., 2016; Prakash et al., 2019).
-Reflection, also referred to as reflective practice (Schön, 1983), or reflective thought, is a multi-faceted concept with multiple meanings (Nguyen et al., 2014; Rogers, 2001; Mann et al., 2009).
-We adopt the definition of the American philosopher John Dewey (1933), who defines reflective thought as the “active, persistent and careful consideration of any belief or supposed form of knowledge in light of the grounds that support it and further conclusion to which it tends” (p9).
-Consequently, reflection can promote reasoning, critical thinking and problem-solving, and thus improve decision-making (Schmidt and Mamede, 2023; Mamede and Schmidt, 2023; Lambe et al., 2016; Prakash et al., 2019; Walger et al., 2016; Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Hess et al., 2015).
+Amongst various forms of increasing cognitive engagement, such as checklists or feedback, stimulating reflection proved to be the most effective (Lambe et al., 2016; Prakash et al., 2019).
 While reflection can occur before, during or after an action (Rogers, 2001), we focus on reflection during decision-making, i.e., reflection-in-action (Schön, 1983).
-In areas with far-reaching decisions, such as law or healthcare, decision-makers have a professional responsibility, such as physicians who are committed to the well-being of patients.
+As mentioned earlier, reflection can promote reasoning, critical thinking and problem-solving, and thus improve decision-making.
+In areas where decisions have far-reaching consequences, such as law or healthcare, decision-makers have a professional responsibility; physicians, for example, are committed to the well-being of patients.
 This professional responsibility involves an epistemological responsibility, which means that decision-makers are responsible for gathering and evaluating relevant information, and knowing the reasons for a particular course of action (van Baalen and Boon, 2015).
-In regard to past actions, knowing and providing reasons for an action allows for backward-looking responsibility or accountability.
-More importantly, however, in regard to future actions, knowing the reasons allows to take forward-looking responsibility by determining and directing actions to achieve a desired state (Loughran, 1996).
-By making it possible to be aware of the reasons for a decision, to formulate and weigh them up accordingly, and to change them if necessary (Mezirow, 1990), reflection in individual cases, i.e., reflection-in-action, can increase epistemological responsibility, both in a backward-looking and, more importantly, in a forward-looking sense.
-Ideally, this forward-looking responsibility leads to fairer and better outcomes for all parties involved (Ueda et al., 2024).
+In regard to past actions, knowing and being able to provide reasons for an action allows for backward-looking responsibility or accountability.
+More importantly, however, knowing the reasons makes it possible to determine and direct future actions in order to achieve a desired state, and thus to take forward-looking responsibility (Loughran, 1996).
+Reflection-in-action can therefore increase professional responsibility, both in a backward-looking and, more importantly, in a forward-looking sense.
+Ideally, this forward-looking responsibility leads to fairer and better decision outcomes for all parties involved (Ueda et al., 2024).
 Reflection-in-action could, for example, prevent a physician from jumping to conclusions after seeing seemingly similar patient cases throughout the day, which could lead to tunnel vision (Mamede and Schmidt, 2023).
-Consequently, to cite Dewey again, reflection “converts action that is merely appetitive, blind and impulsive into intelligent action” (Dewey, 1933, p17).
+Reflection makes it possible to distinguish between relevant and irrelevant information, and, to quote Dewey again, “converts action that is merely appetitive, blind and impulsive into intelligent action” (Dewey, 1933, p.17).
 2.5. The Socratic Questioning Method
-A common technique for facilitating reflection is to ask questions (Osmond and Darlington, 2005).
-One questioning technique that is widely used in education is the Socratic method.
+A common technique to facilitate reflection is to ask questions (Osmond and Darlington, 2005).
+Moreover, questions encourage independent thinking, which can increase confidence in one’s own decision-making, i.e., decision self-efficacy, which in turn can affect the motivation and well-being of the decision-maker (Ryan and Deci, 2000).
+Findings from educational research suggest that motivation functions as an antecedent to cognitive engagement and also increases it (Rotgans and Schmidt, 2011; Blumenfeld et al., 2006; Singh et al., 2022).
+Questions thus seem to promote reflection, which can increase decision self-efficacy and motivation, which in turn can increase cognitive engagement.
+To formulate relevant and effective questions, we want to point out two works, which we also use for our purposes.
+First, Bloom’s taxonomy for educational objectives (Krathwohl, 2002; Anderson and Krathwohl, 2001) categorises cognitive processes into six levels, which are: remembering, understanding, applying, analysing, evaluating, and creating.
+This categorisation is meant to help educators to define learning goals, exercises, and questions accordingly.
+To stimulate reflection, the last three levels, analysing, evaluating, and creating, seem the most promising as they are also components of critical thinking (Paul, 2019), which aims at “judging in a reflective way what to do or what to believe” (Facione, 2000).
+Second, the Socratic method is a questioning method that is widely used in education to develop and promote critical thinking skills (Makhene, 2019; Ho et al., 2023; Dalim, 2022; Torabizadeh et al., 2018).
 The Socratic method consists of systematically asking questions that help to clarify concepts, improve understanding, and uncover gaps in knowledge (Paul, 2019).
-Further, it helps to discover own thoughts, to analyse assumptions, information and inferences, and to arrive at own judgements through own reasoning (Paul, 2019).
-Besides, answering questions promotes reasoning (Alfano, 2019).
-Socratic questions are thus closely linked to critical thinking, which aims at “judging in a reflective way what to do or what to believe” (Facione, 2000).
-In general, independent thinking can promote the decision-autonomy of the decision-maker, which in turn can affect their motivation and well-being (Ryan and Deci, 2000).
-Findings from educational research suggest that motivation increases cognitive engagement (Rotgans and Schmidt, 2011; Blumenfeld et al., 2006) and even functions as a antecedent to it (Singh et al., 2022).
-As soon as students have the opportunity to participate in the organisation of learning activities, their interest and motivation increase (Blumenfeld et al., 2006).
-Without claiming a linear causality, questions seem to encourage reflection, which can increase decision-making autonomy, which can increase motivation, which can increase engagement.
-In view of this, and the connection between critical reflection and reasoning and the associated positive effect on decision-making, we see potential in the transfer of questions, in particular through the Socratic method, from the education sector to the context of machine-supported decision-making.
-The Socratic method is slowly gaining traction in the field of human-machine collaboration, with a conceptual proposal of a virtual assistant that promotes reflection through Socratic questions (Lara and Deckers, 2020), and an educational chatbot based on a large language model and the Socratic method, which was found to improve critical reflection compared to traditional chatbots (Favero et al., 2024).
-There is, however, no systematic approach as to how the Socratic method and questions for critical reflection in general can be transferred to the area of machine-supported decision-making.
-This is partially because the Socratic method is primarily intended for the debate of philosophical issues.
-As such, a current taxonomy for Socratic questions (Paul, 2019, 2010) is primarily aimed at teachers and students to probe and foster thinking.
-We therefore aim to address this gap by transferring this taxonomy for Socratic questions and combining it with other work.
-In doing so, we concretise aforementioned conceptual proposals for human-machine collaborations that go beyond the traditional paradigm of machine recommendations and explanations (Ma et al., 2024b; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023).
+Further, it helps to discover own thoughts, to analyse and evaluate assumptions, information and inferences, and to arrive at judgements through own reasoning (Paul, 2019).
+To help teachers to explore and encourage their students’ thinking through guided questions, a taxonomy of Socratic questions has been developed, defining eight types of questions that probe different parts of thinking, such as assumptions, reasons, or implications (Paul, 2019).
+The Socratic questioning method is slowly gaining traction in the field of human-machine interaction.
+Based on the taxonomy of Socratic questions, a dataset of 110k questions that could be used in counselling and coaching is proposed, including a generation method of prompt-tuning a language model (Ang et al., 2023).
+The resulting questions fall into one of the six categories defined by Paul (2019): 1) clarification (“What do you mean by…?”), 2) probing assumptions (“Why do you assume…?”), 3) probing reasons (“How do you know that…?”), 4) probing implications (“If…, what is likely to happen?”), 5) probing alternatives (“What else should we consider?”), and 6) others that do not conform to the Socratic categories (Ang et al., 2023).
+In addition to this work, a concept of a virtual assistant that fosters moral enhancement through Socratic questions is proposed (Lara and Deckers, 2020), as well as chatbots based on language models that utilise the Socratic method for thought-provoking tutoring (Favero et al., 2024; Liu et al., 2024).
+While these works appear promising, there is currently no systematic approach to using questions to promote critical reflection in decision-making, especially in combination with decision-support systems.
+Given the link between critical reflection and reasoning and the associated positive effect on decision-making, we see potential for the use of questions to promote decision self-efficacy and cognitive engagement.
+We therefore aim to address this gap by transferring an existing taxonomy of Socratic questions (Paul, 2019, 2010) to the context of machine-assisted decision-making.
+In doing so, we also aim for a higher level of granularity as compared to Ang et al.
+(2023), who utilise parts of the Socratic questions taxonomy in unchanged form to generate counselling and coaching questions.
+With our proposed taxonomy we provide a systematic way to formulate relevant questions for critical reflection during decision-making, based on concepts of critical thinking and the Socratic questioning method.
 3. Methodology
-In the following, we will discuss how we have synthesised three earlier works in order to derive our question taxonomy, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revision of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010).
-In the first step, we determine the elements for reflection, i.e., what the question shall relate to.
-To do so, we derive the overall structure of our taxonomy from a taxonomy for Socratic questions (Paul, 2019), which divides questions into the following components of reasoning:
-  • Purpose: Questions that relate to the goal or objective.
-  • Question at issue: Questions that relate to the problem or issue that gave rise to the question.
-  • Information: Questions that relate to background information, such as data, facts, observations or experiences.
-  • Interpretation and Inferences: Questions that relate to how meaning was derived and conclusions drawn.
-  • Concepts: Questions that relate to underlying theories, definitions, or models that define a thought.
-  • Assumptions: Questions that relate to taken for granted presuppositions of a thought.
-  • Implications and Consequences: Questions that focus on implications.
-  • Point of View: Questions that relate to the frame of reference or perspective of a thought.
-We translate these categories and apply them to the context of decision-making in which operators use DSS.
-As such we the derive the following question categories: case information, e.g., input data, (information), relevance of data (information, interpretation and inference), dataset (concepts), causal structure of recommendation (interpretation and inferences), alternatives to recommendation (question, purpose), assumptions and expectations of decision-maker (assumptions, interpretation and inferences), stakeholder preferences (point of view), consequences of recommendation (implications), what-to-be that outcome (purpose), and model behaviour (assumptions) (see column ‘Element for Reflection’ in table 1).
-In the second step, we identify elements, such as data or model behaviour through explanations, that can serve as information for the basis of the content of questions, i.e., how the questions can be enriched.
-To make appropriate use of the information derived from XAI, and given the various techniques available, we turn to a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022).
-The authors cluster more than 50 common questions operators have towards a DSS into the following categories:
+In the following, we will discuss how we constructed our question taxonomy, which we did in two steps: identifying possible questions, and identifying information for formulating questions.
+Overall, we synthesised three prior works, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revision of Bloom’s taxonomy for educational objectives (Krathwohl, 2002; Forehand, 2010).
+The creation of the taxonomy resembles affinity diagramming with a deductive (top-down) approach, for which the Socratic questions functioned as a starting point and the XAI question bank helped us to translate the categories to the context of machine-assisted decision-making (Figure 1).
+3.1. Phase 1: What to question?
+In the first step, we identified relevant aspects of machine-assisted decision-making to which potential questions should relate, i.e.,what elements to reflect on.
+For this purpose, we used a taxonomy for Socratic questions based on critical thinking concepts (Paul, 2019, 2010).
+As Elder and Paul (2010) note:
+“Whenever we think we think for a purpose within a point of view based on assumptions leading to implications and consequences.
+We use data, facts and experiences to make inferences and judgements based on concepts and theories to answer a questions or solve a problem.”
+Accordingly, the authors categorise questions as follows into elements of thought or reasoning that allow to question the structure of thinking (Paul, 2019, 2010):
+  • Purpose: Questions about the agenda, goal or objective.
+  • Question at Issue: Questions about the problem or issue that gave rise to the question.
+  • Information: Questions about background information, such as data, facts, observations or experiences.
+  • Interpretation and Inferences: Questions about how meaning was derived and conclusions drawn.
+  • Concepts: Questions about underlying theories, definitions, or models that define a thought.
+  • Assumptions: Questions about taken for granted presuppositions of a thought.
+  • Implications and Consequences: Questions about implications.
+  • Point of View: Questions about the frame of reference or perspective of a thought.
+While these elements of thought are helpful in general, the provided example questions (Table A1 in the Appendix) cannot be directly applied to the domain of decision-making (with DSSs).
+This is because the taxonomy of Socratic Questions is primarily intended for teachers and students in the classroom.
+We thus took the elements of thought as a starting point for our taxonomy and mapped them to the context of machine-assisted decision-making.
+The element Information, for example, can be mapped, as is also mentioned in the definition, to the dataset the DSS is trained on, as well as data of a specific (patient) case.
+To further specify relevant elements or aspects of machine-assisted decision-making, we used the XAI question bank, in which the authors cluster more than 50 common questions operators have towards a DSS into the following categories (Liao et al., 2020, 2021; Liao and Varshney, 2022):
   • How (global model-wide): Questions about the general logic of the model.
@@
   • Why Not (a different prediction): Questions about the difference to an expected outcome.
-  • How to be That (a different prediction): Questions about how to change
+  • How to be That (a different prediction): Questions about how to change an instance to get a different outcome.
   • How to still be This (the current prediction): Questions about possible changes to still get the same outcome.
@@
   • Output: Questions about how to apply or use the DSS output.
-The authors use these question categories to map them to existing XAI techniques, such as feature importance or counterfactuals.
-This mapping is intended to help designers and developers select the appropriate XAI technique to provide the relevant answer, i.e., explanation.
-In view of the overlap between the categorisation of Socratic questions and the common questions operators have, we use the mapping of questions to XAI techniques provided by the authors of (Liao et al., 2020, 2021; Liao and Varshney, 2022).
-Similarly, we assign XAI techniques to our identified question types in order to enrich possible questions with information (column ‘Useful Information’ in table 1).
-On the one hand, this means that (some) explanations can be (re-)formulated as questions (Danry et al., 2023), e.g., “Does outcome $Y$ follow from feature $x$?”.
-On the other hand, questions can relate to the decisive information extracted from explanations, e.g., “Is feature $x$ the relevant factor to focus on?”.
-So compared to the XAI question bank (Liao et al., 2020, 2021; Liao and Varshney, 2022), we reverse the process by repurposing explanations in order to ask the operators questions that stimulate critical reflection, instead of starting with questions from operators in order to arrive at possible explanations.
-In the final step, and since questions can stimulate different thinking processes, we draw inspiration, albeit only marginally, from a revised version of Bloom’s taxonomy in order to capture the scope of questions (Krathwohl, 2002; Forehand, 2010; Anderson and Krathwohl, 2001).
+Questions from the XAI question bank are about the logic of the model, its performance, the reasons for a prediction, possible changes, the data, and the output.
+This leaves us with, on the one hand, the categorisation of questions based on the structure of thought, and on the other hand, the components of machine-assisted decision-making represented by clusters of common questions in the context of explainable AI (Figure 1).
+Through affinity diagramming we arrived at the following 10 elements: case information, relevance of data, dataset, causal structure of recommendation, alternatives to recommendation, assumptions of decision-maker, stakeholder preferences, consequences of decision, change intervention, and model behaviour.
+Figure 1.
+On the left are the elements of thought based on the taxonomy for Socratic questions (Paul, 2019).
+On the right are the question categories identified in the XAI question bank (Liao et al., 2020).
+Based on these, we define our elements for reflection in machine-assisted decision-making in the middle column.
+To define our final question categories, we followed an iterative process.
+We presented 10 initial categories, including some sample questions, to one orthopaedic surgeon and one senior researcher in orthopaedics who confirmed the completeness of the derived question types.
+We then refined the categories, without adding or removing any, expanded details, added descriptions, and added sample questions to a separate table (Table A3).
+3.2. Phase 2: How to enrich questions?
+After identifying what to question, the next step was to define how to derive and enrich specific questions of said categories.
+For this we again used the XAI question bank.
+For our purposes, however, we reversed the process.
+Instead of starting with questions from operators in order to arrive at suitable explanations, we use XAI techniques to extract information which then inform possible questions (column ‘Useful Information’ in Table 1).
+On the one hand, this means that (some) explanations can be (re-)formulated as questions (Danry et al., 2023), e.g., “Does outcome $Y$ follow from information $x$?”.
+On the other hand, questions can relate to information extracted from explanations, e.g., “Is information $x$ the relevant factor to focus on?”.
+Finally, to capture the scope of questions, we used, albeit only marginally, a revised version of Bloom’s taxonomy (Krathwohl, 2002; Forehand, 2010; Anderson and Krathwohl, 2001).
 In Bloom’s taxonomy, questions are categorised hierarchically into six levels of cognitive processes, which are: remembering, understanding, applying, analysing, evaluating, and creating.
 The last three levels, analysing, evaluating, and creating, are also components of critical thinking (Paul, 2019).
-We thus hypothesise that in order to effectively stimulate critical reflection, questions should address these three higher-level cognitive abilities, where:
+In order to effectively stimulate critical reflection, we thus hypothesise that questions should address these three higher-level cognitive abilities, where:
   • Analysing means to break down information into parts and determining how these parts are connected, and identifying which parts are relevant or irrelevant.
@@
 For example, “What could be done to minimise $Y$?”
-This distinction is meant to help formulate concrete questions for specific use cases.
+This distinction, with the associated verbs, is meant to help formulate concrete questions for specific use cases.
 4. Taxonomy of Questions for Critical Reflection
-In order to help create fruitful questions for critical reflection, we present our taxonomy of question types (table 1).
-We refer to table 2 for some example questions from the medical domain, where a physician uses a DSS in order to find a diagnosis or treatment option for a patient.
-For this, a collection of general clinical questions served as input (Ely, 2000).
+In the following, we will present our taxonomy, which consists of 10 question types or elements for reflection (Table 1).
+In order to demonstrate how to apply our taxonomy, we provide some sample questions from the medical and educational domain in Table A3 and Table A4 respectively.
 Table 1.
-Our question taxonomy, which supports the systematic creation of questions that stimulate critical reflection during decision-making.
-From a taxonomy of Socratic questions we take the elements of thought (Paul, 2019), indicated in italic, and translate them to the decision-making process between human and DSS.
-We provide helpful information that can enrich the questions.
-To map existing XAI techniques to question types, we utilise a question bank for human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022).
+Our proposed taxonomy that supports the systematic creation of questions to stimulate critical reflection in machine-assisted decision-making.
+From a taxonomy of Socratic questions we take the elements of thought (Paul, 2019), indicated in italic, and map them to the context of machine-assisted decision-making.
+To do so and to enrich questions, we utilise and repurpose existing XAI techniques by reversing the mapping between questions and explanations in a question bank for human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022).
 IDs marked with an asterisk (*) indicate questions that address the level of creating, according to Bloom’s categorisation of questions for education (Krathwohl, 2002).
-We refer to table 2 for sample questions.
-ID Element for Reflection (What?) Description (Why?) Useful Information (How?)
+For sample questions, we refer to Table A3 and Table A4 in the Appendix.
+ID Element for Reflection (What?) Description Useful Information (How?)
 Q1 Case Information
-Questions to help further assess, inspect, scrutinise, and contextualise data points, to ensure its quality and reliability. Input Data (e.g., tabular patient data)
+Questions that help further assess, inspect, and contextualise data points to ensure its quality, reliability, and completeness.
+Input Data (e.g., tabular patient data), Datasheets (Gebru et al., 2021)
 Q2 Relevance of Data
-(Information, Interpretation and Inference)
-Questions to help evaluate data to ensure that it is valid by adequately supporting the hypothesis / recommendation.
-Derived explanations (XAI) can be framed as questions.
-Feature Contribution (e.g., SHAP (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016))
+(Information, Interpretation & Inference)
+Questions that help evaluate whether the data adequately supports the hypothesis / recommendation. Feature Contribution (e.g., SHAP (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016))
 Q3 Dataset
-(Concepts )
-Questions to help inspect, assess and judge the assumptions built into the model to ensure that data adequately represents the phenomenon and that the conclusions derived are valid.
-Training Data, Datasheets (Gebru et al., 2021), Model Cards (Mitchell et al., 2019), FactSheets (Arnold et al., 2019), Global Feature Importance (Lundberg et al., 2020)
+Questions that help analyse and assess the data and to ensure it adequately represents the phenomenon (by highlighting limitations or characteristics of the dataset).
+Training Data, Datasheets (Gebru et al., 2021), FactSheets (Arnold et al., 2019)
 Q4 Causal Structure of Recommendation
-(Interpretation and Inference)
+(Interpretation & Inference)
 Questions that help evaluate whether the outcome follows from data to ensure that the causal structure of the model / recommendation is sound.
-Can entail the previous questions of checking whether the data (Q1&Q2) and concept (Q3), i.e., model, are reliable and valid.
-Feature Contribution (e.g., (Lundberg and Lee, 2017), (Ribeiro et al., 2016))
+Feature Contribution (e.g., (Lundberg and Lee, 2017), (Ribeiro et al., 2016)), Counterfactuals
 Q5* Alternatives to Recommendation
 (Question, Purpose)
-Questions that help to consider other possibilities and create other options for actions, to ensure that (larger) solution space is considered.
-Contextual Information, Differential diagnosis, Patient Preferences (Q7)
+Questions that help to consider other possibilities ensuring that the larger solution space is considered.
+Contextual Information (e.g., differential diagnosis or patient preferences (Q7)), Partial Dependence Plot, Counterfactuals,
 Q6 Assumptions and Expectations of Decision-Maker
-Questions that elicit taken for granted assumptions to ensure that the decision-maker is aware of their reasons and can direct their actions accordingly. General Questions
+Questions that elicit taken for granted assumptions to ensure that the decision-maker is aware of their reasons and can direct their actions accordingly.
+General Questions, Confidence of decision-maker
 Q7* Stakeholder Preferences
@@
 Questions that help the decision-maker enquire and take into account the preferences or needs of the people concerned (e.g., patient), in order to ensure a better or broader understanding of the problem at hand.
-General questions
-Q8* Consequences of Recommendation
-Questions that help elicit anticipation and forward-looking responsibility to ensure that (unintended) consequences are considered and mitigated. General questions
-Q9* What to be that Outcome
-Questions to support the exploration of alternative courses of action, by gaining insights from hypothetical scenarios derived from the DSS.
-Feature Perturbation, Counterfactuals (Karimi et al., 2021; Poyiadzi et al., 2020)
+General questions, Preference Elicitation
+Q8* Consequences of Decision
+Questions that help elicit anticipation and forward-looking responsibility to ensure that (unintended) consequences are considered and mitigated.
+Limitations of Model, FactSheets (Arnold et al., 2019), Model Cards (Mitchell et al., 2019)
+Q9* Change Intervention
+Questions that support the exploration of feasible (smaller) interventions to make a desired outcome more likely. Feature Perturbation, Counterfactuals (Karimi et al., 2021; Poyiadzi et al., 2020)
 Q10 Model Behaviour
-Questions that help to assess and evaluate the rules and thresholds of the model.
+(Assumptions, Purpose, Concepts)
+Questions that help to assess and evaluate the assumptions, rules and thresholds built into the model.
 In particular, areas of decision limits, i.e.
 until when does the result remain the same and when does it change.
-Allows to learn about causal structure (Q4).
-Decision Boundaries, Feature Perturbation, Counterfactuals, Model Cards (Mitchell et al., 2019)
-Table 2.
-Based on our question taxonomy in table 1, we provide some sample questions that relate to a medical use case.
+Decision Boundaries, Feature Perturbation, Counterfactuals, Model Cards (Mitchell et al., 2019), Global Feature Importance (Lundberg et al., 2020)
+Q1.
+Questions can relate to the case information provided, which represents information in the form of data.
+Previous operations on the spine, for example, might be indicated by integers, but this value does not specify how long ago the surgery was performed and at which location of the spine it was carried out - both aspects that can influence the effectiveness of further operations.
+Questions can help the decision-maker to inspect and contextualise the data provided to give it semantic depth.
+In addition to ensuring data quality, questions can also relate to the reliability of the data provided to ensure that inconsistencies, such as measurement errors, are ruled out, and that no relevant data is missing.
+As such, Q1 helps the decision-maker get a more holistic picture of the case that goes beyond data-driven abstractions (van Baalen and Boon, 2015).
+Q2.
+Questions can relate to the relevance of data.
+XAI methods make it possible to extract the contribution of the various features to the outcome and rank them accordingly.
+It is possible to use that information to generate an evaluative question, like “How relevant is data $x$?”.
+By questioning the relevance of data points, the decision-maker is encouraged to contextualise the information, e.g., symptom, in a specific case.
+To do so, the decision-maker may wish to check and validate the machine recommendation by other means, e.g., by conducting additional tests.
+So while XAI can show which features have contributed to an outcome, domain knowledge is necessary to evaluate whether these features are important in a particular case.
+Questioning the relevance of the most contributing data point also raises the question whether there are other factors that are overlooked in the decision.
+In addition, considering that different XAI techniques can result in different rankings of the contributing effects of the individual features (Saarela and Jauhiainen, 2021), it is important for decision-makers not to take the extracted information provided by XAI at face value.
+The contradiction between two opposing feature ranks by two different XAI techniques could also lead to a question like “Is $x$ or $z$ more relevant?”.
+Q3.
+Questions can relate to the dataset on which the model was built.
+Cardiovascular diseases, for example, show different symptoms in men and women.
+Most of the collected data that is used to train DSS, however, stems from male patients, which is why DSS often lead to misdiagnosis in women (Li et al., 2023).
+The problem in this case is that certain populations or phenomena are overrepresented in data, which leads to a skewed data distribution (Ueda et al., 2024).
+Besides, data is often only a partial representation of the phenomena under investigation (Broussard, 2019; Binns et al., 2018), as it is impossible to quantify every aspect of the complexity of social life (Birhane, 2021).
+Questions can help to consider limitations of a dataset and take into account information that is not yet included and is (now) known to contribute to the outcome.
+So compared to Q2, which is more concerned with the question of what information is relevant for a single recommendation (local), Q3 asks whether the dataset contains relevant information (global).
+Q4.
+Questions can relate to the (apparent) causal structure between provided case information and resulted machine recommendation.
+Again, based on feature contribution it is possible to ask whether and how the machine recommendation follows from data point $x$, for instance “Does diagnosis $Y$ follow from symptom $x$?”.
+So although this question type merely reformulates a causal explanation (Danry et al., 2023), the aim is to help the decision-maker to analyse and evaluate information.
+Q5.
+Questions can help identify and consider alternatives to the recommendation.
+This calls into question the solution of the DSS and thus, similar to Q4, the appropriateness of a recommendation.
+Different disease have common symptoms, so the physician wants to make sure that another possible diagnosis is not overlooked.
+For this, a differential diagnosis can be helpful (Mamede and Schmidt, 2023).
+At the same time, it can be possible to treat the same condition in different ways.
+Chronic lower back pain, for example, might be treated by surgical or non-surgical interventions (van Hooff et al., 2014), where the feasibility of each option can depend on constraints like available resources or the patient’s preferences (Q7).
+For example, if the purpose of a DSS is to optimise for full recovery, it might recommend surgery that involves a long stay in hospital; another purpose might be to promote the well-being of the patient, who may prefer to spend as little time as possible in hospital, so that the more suitable treatment could be a non-surgical intervention.
+Q5 supports the decision-maker in developing alternative hypotheses and courses of action.
+Q6.
+Questions can relate to the assumptions of the decision-maker.
+On the one hand, these questions can help to elicit tacit knowledge and taken for granted assumptions, which might need to be re-evaluated on a case to case basis.
+This could prevent the decision-maker from jumping to conclusions.
+On the other hand, questions can help the decision-maker to increase their confidence and self-efficacy as they allow them to arrive at their own decision.
+Moreover, assumptions of the decision-maker could also be compared to the machine recommendation.
+Q7.
+Questions can relate to stakeholder preferences.
+Especially in decision contexts in which decision-maker and decision-subject are different persons, such as physician and patient, it is important to take into account different perspectives.
+Ideally, enquiring about the preferences of stakeholders (patient) is part of the decision-maker’s (physician) routine practice, yet these questions can function as a form of safeguard.
+In addition, in case the decision-maker is the decision-subject, questions can also help elicit preferences of the decision-maker themselves.
+Q8.
+Questions can help to anticipate (unintended) consequences of a decision.
+While the decision-maker has ideally weighed up the implications, these questions can promote forward-looking responsibility of the decision-maker.
+Furthermore, with insights from the dataset (Q3) attention could be drawn to certain limitations of a DSS.
+Compared to Q3, the focus of Q8 is on assessing the impact of a recommendation, and supporting the decision-maker in considering appropriate mitigation strategies.
+Q9.
+Questions can relate to change interventions in order to further investigate the space of different alternatives.
+By creating hypothetical scenarios, the feasibility of desired outcomes can be investigated.
+In doing so, the decision-maker is encouraged to consider prior and smaller interventions that could increase the expected effectiveness of a desired treatment.
+To create such question, feature perturbation and counterfactual explanations can be helpful (Karimi et al., 2021; Poyiadzi et al., 2020).
+In contrast to counterfactual explanations, which provide information about what the outcome would have been if the data had been like this, these questions are about whether it is possible to change certain data points in order to make a particular outcome more likely.
+For example, a counterfactual explanation would be “If the salary were higher, the loan would have been accepted”, whereas the question turns into “Is it possible increase the salary, so that the probability of an accepted loan increases?”.
+Although these questions ideally lead to more actionable insights, they might imply unfeasible or unattainable changes and actions, e.g., changing the age of a patient.
+These hypothetical scenarios can nevertheless provide insights into inbuilt associations and decision boundaries of the DSS.
+Q10.
+Questions can relate to the model behaviour.
+For example, if a DSS makes a decision based on age, with a threshold of 50 years, the decision-maker may want to re-evaluate the decision, if the patient is close to that threshold, e.g., 48 years old.
+These questions might also relate to the relevance of data points (Q1), as single data points often function as proxies for other characteristics.
+Age, for example, might be an indication for how well the patient will recover from surgery, yet physical fitness might also play a role.
+This means, focusing only on single data points could be misleading.
+As such, these questions can also give the decision-maker a better understanding of the causal structure built into the model (Q4).
+5. Evaluation of Our Taxonomy
+To evaluate the proposed question taxonomy, we presented a first version with sample questions and the final version in two phases to the same orthopaedic surgeon and orthopaedic researcher in individual meetings.
+Their feedback was positive and they recognised the relevance of the taxonomy.
+As one remarked, given that many question types from the taxonomy are derived from data, it is difficult to ask about data points that are not contained in the dataset.
+[^1] In this case, general questions like “Were all relevant information considered?” could help.
+Nevertheless, they were confident that questions could help to re-evaluate initial assumptions of the physician, and counteract both priming by information from a DSS and habitual priming, e.g., seeing similar yet different patient cases throughout the day.
+Moreover, they mentioned that the questions derived from the taxonomy could help to treat different patient cases more objectively by reducing discrepancies between these cases, such as different patients describing their pain differently, i.e., pain catastrophising.
+5.1. Transferability
+To assess the applicability of our taxonomy in another domain, we discussed the finalised version with two postdoctoral researchers in the field of education and technology, and one doctoral researcher who is also a primary school teacher, in three individual feedback sessions of 30 minutes each.
+They all shared examples of so-called learning analytics dashboards (LAD), with which educators can plan lessons, define learning objectives, track the learning progress of students and create exercises accordingly.
+These LADs also allow teachers to use exercises and learning objectives that the system recommends based on the student’s performance.
+All three researchers therefore confirmed the expected relevance and applicability of our taxonomy for stimulating teachers’ critical reflection in interpreting the student’s progress and integrating recommendations of LADs.
+As one researcher commented: “Teachers feel the need to be more critical, but do not know how”, especially as these LADs can only take part of the student’s trajectory into account while informing and shaping decisions about the student’s future.
+After the three feedback sessions, we created sample questions for the educational domain (Table A4).
+We presented these questions, including our taxonomy, in a lab session attended by 12 people, from doctoral researchers to full professors.
+The research group, which explores the impacts of technologies on learning and teaching practices, recognised the usefulness of our taxonomy.
+As one participant commented, the sample questions we created are very close to the questions teachers already ask themselves.
+Nevertheless, the same researcher continued, our taxonomy could be informative to see which questions are covered and which are left out.
+In addition, another researcher noted that it is very promising that our taxonomy aligns with the questions that teachers ask in relation to algorithmic learning goal recommendations.
+In comparison to the medical sample questions (Table A3), the wording of the educational sample questions differs.
+The structure and elements of critical reflection however remain the same for both.
+We therefore maintain that our taxonomy is transferable and can be applied to other decision-making domains without adaptation as it builds on general concepts of critical thinking.
+5.2. Effectiveness of Questions
+Another type of evaluation, beyond the scope of this paper, is to assess the effectiveness of the questions created from the taxonomy in decision-making, and the extent to which they promote critical reflection and cognitive engagement of the decision-maker.
+Existing studies, primarily from the educational domain, point to the positive effects of Socratic questioning strategies on moral reasoning in nursing (Torabizadeh et al., 2018) and critical thinking skills during (healthcare) education of students (Ishtiaq et al., 2024; Jameel et al., 2025; Yang et al., 2005; Makhene, 2019; Ho et al., 2023).
+These studies demonstrate the validity and robustness of the Socratic questioning approach on which our proposed taxonomy is based, which makes the prospects of our taxonomy seem promising.
+To measure reflection and cognitive engagement, it is important to consider the entire decision-making process, including the reasoning process, whether alternatives were considered, and how confident the decision-maker is in the final decision.
+Current evaluation methods for decision-making and overreliance on DSS, however, typically only focus on the outcome of a decision, i.e., decision accuracy (Guo et al., 2024; Klingbeil et al., 2024).
+As such, evaluating the effects of questions in promoting critical reflection and decision-making remains an open challenge.
+Existing methods for measuring cognitive engagement are associated with considerable effort and technical barriers, such as electroencephalography (EEG) (Hassib et al., 2017; Berka et al., 2007), or the measurement of pupil dilation (Van Der Wel and Van Steenbergen, 2018).
+Consequently, an evaluation framework would need to be developed in future work, e.g., a self-report scale that takes the entire decision-making process into account.
+For this, inspiration can be drawn from existing scales for factors associated with cognitive engagement, such as self-efficacy, motivation, and confidence (Lee et al., 2025; Hart et al., 2014; Greene, 2015), as well as from methods in the educational domain for assessing student’s cognitive engagement with technology (Vongkulluksn et al., 2022).
+6. Discussion
+We have presented a taxonomy of question types to support the systematic creation of questions that stimulate critical reflection during machine-assisted decision-making.
+To do so, we combined and utilised a taxonomy of Socratic questions and a question bank for explainable AI.
+We do not claim that there is a direct mapping between the taxonomy of Socratic questions and the question categories from the XAI question bank.
+Nevertheless, as Figure 1 shows, the questions identified for the human-centred XAI approach are very model- and data-centric.
+In that regard, and referring back to the need for explanations beyond the model (section 2.2), our taxonomy can help to take into account human factors like assumptions of the decision-maker and stakeholder preferences, as well as consequences of a decision.
+In doing so, questions can stimulate discussion about the machine recommendation and how it can be integrated into the broader decision-making process.
+6.1. Different Stakeholders
+Our taxonomy is primarily targeted at decision-makers who operate DSS to make informed decisions.
+Consequently, the aim of our taxonomy is not to formulate questions for the decision-maker who then evaluates the performance of the DSS.
+This would unfairly shift the responsibility from developers and regulators to operators.
+The deployed models must be evaluated and tested beforehand, especially in crucial settings such as healthcare, which requires a shift from individual human oversight to institutional oversight (Green, 2022).
+Considering this, other stakeholders could also benefit from critical reflection and thus our taxonomy.
+Developers, for example, might ask themselves which taken for granted assumptions they have while constructing a model and whether the used dataset is adequate to solve a given problem.
+Auditors assessing compliance could benefit from our taxonomy when defining objectives for an audit of the model or dataset (Raji et al., 2020; Birhane et al., 2024).
+Moreover, the taxonomy could be used as a form of supporting AI literacy to help people remain more critical of machine-generated information, especially in view of misinformation (Zavolokina et al., 2024).
+Patients could be supported in identifying (treatment) preferences, and asking relevant questions to their healthcare providers (operating DSS), or while interacting with a medical chatbot themselves.
+Students could be taught to question the relevance of retrieved information and could be helped to remain cognitively engaged in their technology-aided learning.
+Our proposed taxonomy could therefore benefit various stakeholders involved in the design, development, testing, and use of DSS or machine learning systems.
+6.2. Creation of Questions
+For questions to be meaningful and to support reflection effectively, researchers or designers ideally create them with input and feedback from domain experts, e.g., physicians.
+While creating the sample questions for the educational domain (Table A4) we realised that it is difficult to create questions without having much knowledge about the respective use case, access to data, and the functionality of the DSS or learning analytics dashboard (LAD).
+Nevertheless, uncovering these blind spots might enable further inquiry into how the system works, and open up a dialogue between developers or service providers and domain experts or affected organisations, such as hospitals or schools.
+Furthermore, certain questions might fall into multiple categories of our taxonomy, as different question types can overlap or build on other categories, such as the causal structure of a recommendation (Q4) that can entail questions about the data (Q1&Q3).
+We believe, however, that this is not a major problem and could rather help to formulate questions more precisely so that they can be delineated.
+The overlap also shows how different elements or categories are connected to each other.
+To help formulate questions, we have presented Bloom’s taxonomy.
+We argue that questions should address the higher levels of critical thinking, that is evaluating, analysing, and creating.
+As such, it might be that questions cannot be answered directly and lead to further questions (Paul, 2019).
+To motivate further inquiry and engagement of the decision-maker, it could be helpful, as mentioned by the orthopaedic surgeon we spoke to, to provide a brief explanation why a question should be entertained.
+For example, the statistical effect of the patient’s age on the outcome, according to the DSS, could be indicated in the question about the age.
+In view of this, however, it must also be considered that questions can be leading and point in certain directions, which might lead the decision-maker to make an undesirable decision.
+It is thus again important that domain experts are involved in the formulation of questions.
+6.3. Presentation of Questions
+Regarding the presentation of questions to decision-makers, one of the educational researchers we spoke to suggested that a poster of our taxonomy with sample questions could be displayed in the staff room to sensitise teachers to the different elements of the machine-assisted decision-making process.
+Alongside an analogue form and building on the above-mentioned proposals of DSS supporting deliberation and reflection (section 2.3), our taxonomy could inform the design of a technical implementation.
+As such, we envision a reflection machine (RM) that functions as an additional system or component of a DSS.
+The RM can take into account the input data (e.g., patient case information), the functioning of the DSS, e.g., through XAI methods such as feature importance, as well as the recommendations.
+With this information and the use of our taxonomy relevant questions can be generated and presented to the decision-maker.
+We advise against the use of language models, as the generated questions could be irrelevant and unreliable (Weidinger et al., 2022) and chatbots might lead to further overreliance (Anderl et al., 2024; Maeda and Quan-Haase, 2024).
+Besides, supporting the decision-making process with questions might be an unusual and unlearned interaction.
+Hence, evaluating the acceptance and usability (Lund, 2001; Venkatesh et al., 2016) of such an RM with specific questions is another consideration while assessing the effectiveness of questions.
+6.4. Cognitive Load
+While future work needs to evaluate the effectiveness of questions, it could be that questions interrupt the workflow and increase the cognitive load of the decision-maker.
+Similarly, in the aforementioned study, where explanations were presented at different times as a form of cognitive intervention, participants reported an increased cognitive load from these interventions, which is why they favoured them less (Buçinca et al., 2021).
+Ultimately, the cognitive burden, however, is unavoidable, and inherent to critical reflection.
+As John Dewey (1933) mentions on the difficulty of reflective thinking:
+[…] yet thinking need not be reflective.
+For the person may not be sufficiently critical about the ideas that occur to him.
+He may jump at a conclusion without weighing the grounds on which it rests; he may forego or unduly shorten the act of hunting, inquiring; he may take the first ‘answer’ or solution, that comes to him because of mental sloth, torpor, impatience to get something settled.
+One can think reflectively only when one is willing to endure suspense and to undergo the trouble of searching (p.17).
+The aim of increasing cognitive engagement therefore leads to an increase in cognitive load.
+Nevertheless, the right amount of effort required and effectiveness needs to be found.
+Questions must be well integrated into decision-making processes in order to enable the operator to constructively take the posed question into account.
+This also depends at least in part on the competences, decision-making style, and need for cognition of the decision-maker.
+7. Conclusion
+We have provided a taxonomy of 10 question types aimed at promoting critical reflection in machine-assisted decision-making.
+Careful selection and presentation of these question types in the right circumstances (e.g., when time for reflection is available) can increase cognitive engagement and support the effective human oversight required by the European AI Act.
+Acknowledgements.
+We thank Sebastiaan van de Groes and Miranda van Hooff, both at Radboud UMC, for their input on the construction of the taxonomy.
+Many thanks to Josh Ring and the Adaptive Learning Lab, in particular to Susan Verbunt-Janssen, Susanne de Mooij, and Zowi Vermeire for their valuable insights, and helping with the sample questions for the educational domain.
+Thanks to Cătălina Rățală and the Decision Science Group for their questions and feedback on a presentation of our paper.
+Comments from three anonymous ACM FAccT reviewers on an earlier version helped us to improve the paper.
+This research is funded by the Donders Centre for Cognition.
+Appendix A Sample Questions
+A.1. The Socratic Question Taxonomy
+Table A1. The following questions are taken from the taxonomy of Socratic Questions (Paul, 2019). The authors provide a more comprehensive list of questions.
+Element of Thought Sample Questions
+Purpose What are we trying to accomplish? What is the purpose of …?
+Question Is this question the best one to focus on? How can we find out?
+Information On what information are you basing that comment? Are those reasons adequate?
+Interpretation & Inference How did you reach that conclusion? How shall we interpret these data?
+Concepts What is the main idea you are using in your reasoning? What main theories do we need to consider in figuring out ….?
+Assumptions What are you taking for granted here? Why do you think the assumption holds here?
+Implications If we do this, what is likely to happen as a result? What are you implying by that?
+Viewpoints From what point of view are you looking at this? What would someone who disagrees say?
+A.2. The XAI Question Bank
+Table A2. The following questions are taken from the XAI question bank (Liao et al., 2021). The authors provide a more comprehensive list of questions.
+Element Sample Questions
+Data What kind of data was the system trained on? What is the sample size of the training data?
+Output What kind of output does the system give? How should the output fit in my workflow?
+Performance How accurate/precise/reliable are the predictions? How often does the system make mistakes?
+How (global) How does the system make predictions? What kind of algorithm is used?
+Why Why/how is this instance given this prediction? What feature(s) of this instance determine the system’s prediction of it?
+Why not Why is this instance NOT predicted to be [a different outcome Q]? Why is this instance predicted [P instead of a different outcome Q]?
+What if What would the system predict if this instance changes to….? What would the system predict for [a different instance]?
+How to be that How should this instance change to get a different prediction Q? What is the minimum change required for this instance to get a different prediction Q?
+How to still be this What is the scope of change permitted for this instance to still get the same prediction? What kind of instance gets the same prediction?
+Others How/why will the system change/adapt/improve/drift over time? Can I, and if so, how do I, improve the system?
+A.3. Our Taxonomy applied to Healthcare
+Table A3.
+Based on our question taxonomy in Table 1, we provide some sample questions that relate to the medical domain, where a physician uses a decision-support system (DSS) in order to find a diagnosis or treatment option for a patient.
 A collection of general clinical questions serves as input (Ely, 2000).
-Relevant questions will vary from domain to domain, these questions should thus be taken as suggestions.
+Relevant questions will vary depending on the use case, these questions should thus be taken as inspiration.
 ID Element for Reflection Sample Questions
-Q1 Case Information Is the provided data complete?
-How do you think the patient understood the self-assessment questionnaire?
-When was the indicated surgery performed?
+Q1 Case Information How does the patient compare against the data distribution?
+Does the data adequately represent the patient’s situation?
+How do you think the patient understood the self-assessment questionnaire that functions as input to the DSS?
+When was the indicated surgery performed (previous surgeries decrease the effectiveness of future surgery)?
 Why did provider $x$ treat the patient this way?
@@
 What is the likelihood that symptom $x$ is coming from condition $y$?
-Q3 Dataset Does the data adequately represent the phenomena under investigation?
-Is the model up-to-date?
+Does feature $x$ [salary] that is part of the data affect the decision?
+Q3 Dataset Does the dataset adequately represent the phenomena under investigation?
+Is the data up-to-date?
 Is there relevant data missing in the dataset?
-Did you consider $x$, which the model does not consider?
+Did you consider the patient’s BMI measure, which the model does not consider, but which influences the result?
+Where was the data collected?
+What is the data sample size?
+Is limitation $z$ of the dataset important to consider (information ideally provided by Datasheet)?
+Is it important that the data was collected more than 5 years ago?
 Q4 Causal Structure of Recommendation What is the likelihood that this patient has condition $y$ (given findings $x_{1}$, $x_{2}$,…, $x_{n}$)?
-Which information supports/contradicts diagnosis?
+Does diagnosis $y$ follow from symptom $x$?
+Which information supports/contradicts the diagnosis?
 How good is test $x$ in situation $y$?
@@
 How do you distinguish between conditions $y_{1}$, $y_{2}$?
-Could this patient have condition y (given findings $x_{1}$, $x_{2}$)?
+Could this patient have condition $z$ (given findings $x_{1}$, $x_{2}$)?
 Q6 Assumptions and Expectations of Decision-Maker How does the machine recommendation compare to your assumptions?
@@
 Are there alternative assumptions you could make?
-Q7 Stakeholder Preferences Does the patient have any preferences that might require a procedure that differs from the recommendation? Does the patient have strong preferences for treatment $y$?
-Q8 Consequences of Recommendation Are there any unintended consequences of treatment $y$?
+What symptoms are you basing your decision on?
+Q7 Stakeholder Preferences Does the patient have any preferences that might require a procedure that differs from the recommendation?
+Does the patient have strong preferences for treatment $y$?
+Are there certain situational circumstances that prevent the patient from recovering from surgery for several weeks?
+Q8 Consequences of Decision Are there any unintended consequences of treatment $y$?
 What are the ethical/legal considerations in situation $y$?
 What are the administrative considerations in situation y?
-Q9 What to be that Outcome Is it possible to change the patient’s expectations to increase the likelihood of effective surgery?
-Q10 Model Behaviour Would you suggest the same treatment if the patient were 5 years older?
-Q1.
-Questions can address available information, such as data in the form of provided case information.
-Previous operations on the spine, for example, might be indicated by integers, but this value does not specify how long ago the surgery was performed and at which location of the spine it was carried out - both aspects that can influence the effectiveness of further operations.
-Questions can help the decision-maker to inspect and contextualise the data provided to give it semantic depth.
-In addition to ensuring the quality of the data, question can relate to the reliability of the data provided to ensure that inconsistencies, such as measurement errors, are ruled out, or that the data adequately reflect the intended purpose, such as the pain scores self-assessed by patients using a questionnaire.
-Q2.
-The decisive factors for a specific machine recommendation, i.e., relevance of single data points, are another piece of information that can be questioned.
-XAI methods, such as SHAP (Lundberg and Lee, 2017) or LIME (Ribeiro et al., 2016), make it possible to extract the contribution of the various features to the outcome and rank them accordingly.
-In order to stimulate critical reflection, it is possible to use that information to generate an evaluative question, like “Is data point $x$ the one to focus on?” or “How relevant is data $x$?”.
-By questioning the relevance of decisive data points, the decision-maker is encouraged to contextualise the information, e.g., symptom, in a specific case.
-To do so, the decision-maker may wish to check and validate the machine recommendation by other means, e.g., by conducting additional tests.
-So while XAI can show which features have contributed to an outcome, domain knowledge is necessary to evaluate whether these features are important.
-Questioning the relevance of the most contributing data point also raises the question whether there are other factors that are overlooked in the decision.
-In addition, considering that different XAI techniques can result in different rankings of the contributing effects of the individual features (Saarela and Jauhiainen, 2021), it is important for decision-makers not to take the extracted information provided by XAI at face value.
-The contradiction between two opposing feature ranks by two different XAI techniques could also lead to a question like “Is $x$ or $z$ more relevant?”.
-Consequently, questions of this type (Q2) might overlap with Q4 that relates to the causal structure of recommendations.
-Q3.
-Questions can address the concepts of an approach or problematisation, which in the context of joint decision-making translates to the dataset on which the model is built.
-Cardiovascular diseases, for example, show different symptoms in men and women.
-Most of the collected data that is used to train DSS, however, stems from male patients, which is why DSS often lead to misdiagnosis in women (Li et al., 2023).
-The problem in this case is that certain populations or phenomena are overrepresented in data, which leads to a skewed data distribution (Ueda et al., 2024).
-Besides, data is often only a partial representation of the phenomena under investigation (Broussard, 2019; Binns et al., 2018), as it is impossible to quantify every aspect of the complexity of social life (Birhane, 2021).
-Data might have been left out at training the model for different reasons, or was not available in the first place.
-Consequently, questions like “Does the data adequately represent the phenomena under investigation?”, or “Is more evidence necessary?”, can help to validate the functioning and appropriateness of the model and to learn about its limitations.
-In addition, questions like “Have you considered information x?” could be helpful if certain data is not included in the dataset and is therefore not considered by the model, although it is (now) known that the missing information contributes to the outcome.
-So compared to Q2, which is more concerned with the question of what information is relevant in an existing dataset, Q3 asks whether the dataset itself contains the relevant information.
-To create such questions, documentation of the dataset can be very useful (Gebru et al., 2021; Arnold et al., 2019).
-Q4.
-To further assess the interpretation and inferences of the DSS, questions can help to investigate the (apparent) causal structure of provided case information and resulted machine recommendation.
-Again, based on feature contribution it is possible to ask whether and how the machine recommendation follows from data point $x$, for instance “Does diagnosis $Y$ follow from symptom $x$?”.
-So although this question type ‘merely’ reformulates a causal explanation, the aim is to help the decision-maker to analyse and evaluate information.
-Q5.
-Questions can relate to the question or problem to which a DSS is trying to provide an answer, i.e., the purpose of the DSS.
-So similar to Q4 these questions can address the appropriateness of a recommendation.
-The focus of Q5, however, is on helping the decision-maker to consider alternatives to the recommendation.
-Different disease have common symptoms, so the physician wants to make sure that another possible diagnosis is not overlooked.
-For this, a differential diagnosis can be helpful (Mamede and Schmidt, 2023).
-At the same time, it can be possible to treat the same condition in different ways.
-Chronic lower back pain, for example, might be treated by surgical or non-surgical interventions (van Hooff et al., 2014), where the feasibility of each option can depend on constraints like available resources or the patient’s preferences (Q7).
-For example, if the purpose of a DSS is to optimise for full recovery, it might recommend surgery that involves a long stay in hospital; another purpose might be to promote the well-being of the patient, who may prefer to spend as little time as possible in hospital, so that the more suitable treatment could be a combined physical and psychological program.
-In this way, there is an overlap with question type Q7 that relates to stakeholder preferences.
-Nevertheless, the focus of Q5 is on whether alternative recommendations should be considered, and on supporting the decision-maker in developing alternative hypotheses and courses of action.
-Q6.
-Questions can address assumptions and expectations of the decision-maker.
-On the one hand, these questions can help to elicit tacit knowledge and taken for granted assumptions, which might need to be re-evaluated on a case to case basis.
-This could prevent the decision-maker from jumping to conclusions.
-On the other hand, questions that address the expectations of the decision-maker can increase their confidence and allow them to arrive at their own decision.
-In other words, these questions can help to increase the decision-autonomy of the operator.
-Example questions could be, “Does the recommendation match your assumptions?
-If so, why (not)?”.
-General questions allow to take into account domain expertise that can address factors which might not be available in quantified data (Q3).
-As such, questions could allow the decision-maker to obtain a more holistic picture of the problem that goes beyond static rules implemented in DSS (van Baalen and Boon, 2015) (Q5).
-Q7.
-Questions can relate to different perspectives or points of view.
-Especially in decision context in which decision-maker and decision-subject are different persons, such as physician and patient, it is important to take into account stakeholder preferences.
-As such, questions like “Have you considered the preferences of the patient?”, or “Are there certain situational circumstances that prevent the patient from recovering from surgery for several weeks?”, can help to ensure the decision-maker considers other parties involved and treat the patient in a more personalised way.
-Ideally, enquiring about the patient’s preferences is part of the physician’s routine practice during consultation, yet these questions can function as a form of safeguard.
-Q8.
-Questions can help to anticipate (unintended) consequences of a decision.
-These questions can elicit forward-looking responsibility of the decision-maker, for example, by asking “Have you considered the implications of administering drug $x$?” While the decision-maker has ideally weighed up the implications, these questions can function as a reminder.
-Furthermore, with insights from the dataset (Q3) attention could be drawn to certain limitations of a DSS.
-Compared to Q3, the focus of Q8 is on assessing the impact of a recommendation that may be based on underrepresented data, and supporting the decision-maker in considering appropriate mitigation strategies.
-Q9.
-Questions can take the form of what-to-be that outcome in order to further investigate the space of different alternatives.
-By creating hypothetical scenarios, the feasibility of desired outcomes can be investigated.
-So for example, ‘Is it possible to lower level $z$, this could probably make treatment $y$ more effective?” In doing so, the decision-maker is encouraged to consider prior and smaller interventions that could increase the expected effectiveness of a desired treatment.
-To create such question, feature perturbation and counterfactual explanations are helpful (Karimi et al., 2021; Poyiadzi et al., 2020).
-In contrast to counterfactual explanations, which provide information about what the outcome would have been if the data had been like this, these questions are about whether it is possible to change certain data points in order to make a particular outcome more likely.
-For example, a counterfactual explanation would be “If finding $x$ were not present, the condition would be $y$”, whereas the question turns into “Is it possible to lower $x$, so that the likelihood of condition $y$ increases?”.
-Although these questions ideally lead to more actionable insights, they might imply unfeasible or unattainable changes and actions, e.g., changing the age of a patient.
-These hypothetical scenarios can nevertheless provide insights into inbuilt associations and thresholds of the DSS decision-boundaries.
-Q10.
-Questions can address assumptions, as mentions before.
-In addition to the assumptions of the decision-maker, questions can refer to built-in assumptions or rules of model behaviour.
-For example, if a DSS makes a decision based on age, with a threshold of 50 years, the decision-maker may want to re-evaluate the decision, if the patient is close to that threshold, e.g., 48 years old.
-An example question could be “Would you recommend the same treatment if the patient were 3 years older, which might decrease its effectiveness?”.
-These questions might also relate to the relevance of data points (Q1), as single data points often function as proxies for other characteristics.
-Age, for example, might be an indication for how well the patient will recover from surgery, yet physical fitness might also play a role.
-This means, focusing only on single data points could be misleading.
-As such, these questions can also give the decision-maker a better understanding of the causal structure built into the model (Q4).
-Documentation on model behaviour can be a source from which questions can be derived (Mitchell et al., 2019).
-5. Discussion, Future Work, and Concluding Remarks
-We have proposed a taxonomy of question types to support the systematic creation of questions that can stimulate critical reflection during machine-supported decision-making.
-In doing so, we have focused on the medical domain by providing some example questions.
-We maintain that our taxonomy is transferable to other decision-making domains.
-Whilst the wording of the questions may change when applied to other contexts, the general structure and elements of critical reflection remain the same.
-Future work is required to test and validate the generalisability of our taxonomy.
-Next to that some considerations remain.
-The tone of questions is important.
-For natural language explanations, it was found that the tone and assertiveness has an effect on decision-making (Calisto et al., 2023).
-Although questions are in general less assertive compared to explanations, it could be assumed that the tone of questions has a similar effect on how or whether the decision-maker engages with the questions.
-A question could be framed in a more imposing way, such as, “Can you check for symptom $x$?”, rather than in a suggestive way, e.g., “Have you checked for other symptoms?”.
-To examine the right tone, we have presented Bloom’s taxonomy.
-We argue that questions should follow the higher levels of critical thinking, that is evaluating, analysing, and creating.
-As such, it is likely to assume that the questions should be phrased in an open-ended manner.
-Simple yes-or-no questions could be overlooked or act as checkboxes without eliciting cognitive engagement.
-Besides, it is not necessarily the case that definite answers have to be found to questions.
-Rather, the posed question can lead to further questions and thus inquiry of the decision-maker (Paul, 2019).
-In order to support reflection effectively, it is therefore necessary to create questions in a controlled and systematic way.
-Consequently, it might be advisable to (currently) refrain from using large language models (LLMs), as the output risks being irrelevant and unreliable (Weidinger et al., 2022).
-Besides, conversational presentation of information in the form of chatbots might lead to further overreliance (Anderl et al., 2024), as humanised or anthropomorphised chatbots can lead to an illusion of reciprocity and care (Maeda and Quan-Haase, 2024).
-Future work could nevertheless investigate how to stimulate continuous reflection by processing the answers of the operator and asking follow-up questions.
-Ideally, an interactive interface would allow the decision-maker to modify data points accordingly.
-At the same time, however, decision-makers should not be burdened with too many or too frequent questions.
-In the aforementioned study on cognitive interventions, the cognitive load of these interventions on the decision-maker was reported, which is why participants favoured them less (Buçinca et al., 2021).
-The right balance of attention and engagement must be found.
-The operator should be enabled to constructively take the posed question into account, which also depends at least in part on the competences and confidence of the decision-maker.
-Ultimately, the cognitive burden is unavoidable, and inherent to critical reflection.
-As John Dewey (1933) mentions on the difficulty of reflective thinking:
-yet thinking need not be reflective.
-For the person may not be sufficiently critical about the ideas that occur to him.
-He may jump at a conclusion without weighing the grounds on which it rests; he may forego or unduly shorten the act of hunting, inquiring; he may take the first ‘answer’ or solution, that comes to him because of mental sloth, torpor, impatience to get something settled.
-One can think reflectively only when one is willing to endure suspense and to undergo the trouble of searching (p17).
-With this in mind, and given that questions during the decision-making process are likely to reduce efficiency by interrupting the workflow and taking up more time, the design and implementation of a potential ‘reflection machine’ that supports reflection seems less desirable.
-However, in domains with high-impact decisions, such as healthcare, it should be justified to make a deliberate and thus perhaps more time-consuming decision, as the outcome can have life-changing effects.
-Considering that DSS recycle past assumptions with potentially harmful consequences (Burrell, 2024), critical reflection becomes increasingly important to promote responsible decision-making.
-In particular, forward-looking responsibility, which can be strengthened through reflection, ideally leads to a more open and just future.
-In this paper we provided a taxonomy of questions aimed at promoting critical reflection on machine recommendations.
-Careful selection and presentation of these question types in the right circumstances (e.g., when time for reflection is available), and appropriate combination with recommendations and explanations can support the effective human oversight required by the European AI Act.
+What were the consequences of treatment $y$ in previous similar patient cases?
+What are the implications of administering drug $x$?
+Q9 Change Intervention Is it possible to reduce $x$, so that $y$ becomes more likely? Is it possible to change the patient’s expectations to increase the likelihood of effective surgery?
+Q10 Model Behaviour Would you suggest the same treatment if the patient were 5 years older? Did you consider that the model makes errors in 1 out of 10 cases?
+A.4. Our Taxonomy applied to Education
+Table A4.
+Based on our question taxonomy in Table 1, we provide some sample questions that relate to an educational use case where teachers use learning analytics dashboards to plan lessons, track student progress, set learning goals, and create exercises.
+The following questions were created with the help of three researchers in education and technology, one of whom is a teacher.
+Relevant questions will vary depending on use case, these questions should thus be taken as inspiration.
+ID Element for Reflection Sample Questions
+Q1 Case Information Does the data adequately represent the student’s progress? Are there other factors outside the LAD that affect the student’s progress?
+Q2 Relevance of Data Does the failed exercise represent the student’s overall performance?
+Is the recommended exercise relevant to the learning goal?
+How important is the result of the failed / passed exercise for the overall progress?
+Q3 Dataset What (student) data was the model trained on? Which metrics are included in the dataset?
+Q4 Causal Structure of Recommendation Does the recommended exercise contribute to the learning goal? How effective is the recommended exercise for the learning goal?
+Q5 Alternatives to Recommendation Are there other suitable exercises that the students can do without their devices (what is pedagogically needed for my class)?
+What other exercise can be done to achieve the learning goal?
+Should the student work on a learning goal based on their own progress or based on the level of the whole class?
+Q6 Assumptions and Expectations of Decision-Maker What are my assumptions about the student? Are my assumptions about the student adequate?
+Q7 Stakeholder Preferences Does the student have any special needs?
+Can the student benefit from other exercises / help?
+Are there any (extraordinary) circumstances of the student that should be considered?
+Q8 Consequences of Decision Does the recommended exercise contribute to the learning goal?
+What does it do to the emotional well-being of my student if they have to repeat exercises multiple times based on LAD recommendations?
+What are the downsides of this exercise?
+What other learning objectives are being used in this exercise that the student could be struggling with?
+Q9 Change Intervention Is it possible to increase the student’s motivation to make the learning goal more likely?
+Q10 Model Behaviour Does the LAD contain learning theory / methodology X / adhere to our school-wide policy / goal?
+Is the model’s assumption that the student will not achieve the learning goal if fail two exercises appropriate?